### v1

# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

Abstract
We investigate a new diffuse-interface model that describes creeping two-phase flows (i.e., flows exhibiting a low Reynolds number), especially flows that permeate a porous medium. The system of equations consists of a Brinkman equation for the volume averaged velocity field as well as a convective Cahn–Hilliard equation with dynamic boundary conditions for the phase-field, which describes the location of the two fluids within the domain. The dynamic boundary conditions are incorporated to model the interaction of the fluids with the wall of the container more precisely. In particular, they allow for a dynamic evolution of the contact angle between the interface separating the fluids and the boundary, and also for a convection-induced motion of the corresponding contact line. For our model, we first prove the existence of global-in-time weak solutions in the case where regular potentials are used in the Cahn–Hilliard subsystem. In this case, we can further show the uniqueness of the weak solution under suitable additional assumptions. Moreover, we further prove the existence of weak solutions in the case of singular potentials. Therefore, we regularize such singular potentials by a Yosida approximation, such that the results for regular potentials can be applied, and eventually pass to the limit in this approximation scheme.
Keywords: two-phase flows, porous media, Cahn–Hilliard equation, Brinkman equation, dynamic boundary conditions, bulk-surface interaction.
Mathematics Subject Classification: 35D30, 35K35, 35K86, 35B65, 76D03, 76T06
1   Introduction
The mathematical study of two-phase flows is an important topic in many areas of applied science such as engineering, chemistry and biology. To predict the motion of two fluids, it is crucial to understand how the interface between these fluids evolves. To provide a mathematical description of this interface, two fundamental methods have been developed: the sharp-interface approach and the diffuse-interface approach. In the former, the interface is represented as a hypersurface evolving in the surrounding domain. The occurring quantities (e.g., the velocity fields) are then described by a free boundary problem. In the latter, the fluids are represented by a phase-field function which is expected to attain values close to $1$ in the region where the first fluid is present, and close to $-1$ in the region where the second fluid is located. However, unlike in sharp-interface models, this phase-field does not jump between the values $1$ and $-1$ but exhibits a continuous transition between these values in a thin tubular neighborhood around the boundary between the fluids. This tubular neighborhood is referred to as the diffuse interface and its thickness is proportional to a small parameter $\epsilon>0$. For a comparison of sharp-interface methods and diffuse-interface methods, we refer to [28, 4, 35, 55]. We point out that, even though the sharp-interface and the diffuse-interface approach are conceptually different, they can, in general, be related by the sharp-interface limit in which a parameter related to the thickness of the diffuse interface is sent to zero.
In the context of diffuse-interface models, such models in which the phase-field is described by a Cahn–Hilliard type equation have become particularly popular. One of the most widely used models for describing the motion of two viscous, incompressible fluids with matched (constant) densities is the Model H. It was first proposed in [45] and was later rigorously derived in [43]. The PDE system consists of an incompressible Navier–Stokes equation coupled with a convective Cahn–Hilliard equation. In terms of mathematical analysis, the Model H was investigated quite extensively, see, e.g., in [1, 10, 30, 41]. Further generalizations of this model can be found in [51, 11, 27, 57, 58, 35, 44, 29].
One drawback of the Model H is that it can merely be used to describe the situation in which the fluids have the same individual density. To overcome this issue, a thermodynamically consistent diffuse-interface model for incompressible two-phase flows with possibly unmatched densities was derived in the seminal work [6]. This model is usually referred to as the AGG model. Concerning mathematical analysis of this model, we refer the reader to [2, 3, 7, 39, 38, 5]. The connection between the AGG model and the two-phase Navier–Stokes free boundary problem is explained in [6, 4].
Even though the AGG model and the Model H subject to the classical boundary conditions (i.e., a no-slip boundary condition for the velocity field and homogeneous Neumann boundary conditions for the convective Cahn–Hilliard equation) are well suited to describe the motion of the fluids in the the interior of the considered domain, they still inherit some limitations from the underlying (convective) Cahn–Hilliard system with homogeneous Neumann boundary conditions. The main limitations are:
  1. (L1)
    The homogeneous Neumann condition on the phase-field enforces that the diffuse interface always intersects the boundary at a perfect ninety degree contact angle. This will not be fulfilled in many applications. In general, the contact angle might even change dynamically over the course of time.
  2. (L2)
    The no-slip boundary condition on the velocity field makes the model not very suitable to describe general moving contact line phenomena. As the trace of the velocity field at the boundary is fixed to be identically zero, any motion of the contact line of the diffuse interface can be caused only by diffusive but not by convective effects.
  3. (L3)
    The mass of the fluids in the bulk is conserved. Therefore, a transfer of material between the bulk and the boundary (caused, e.g., by absorption processes or chemical reactions) cannot be described.
A more detailed discussion can be found in [40].
To overcome the aforementioned restrictions (L1) and (L2), a class of dynamic boundary conditions was derived in [56]. It involves an Allen–Cahn type dynamic boundary condition for the phase-field coupled to a generalized Navier-slip boundary condition for the velocity field. The Model H subject to this boundary condition was analyzed in [31] whereas the AGG model subject to this boundary condition was investigated in [32].
Recently, a thermodynamically consistent generalization of the AGG model subject to another class of dynamic boundary conditions was derived in [40]. Here, the boundary condition consists of a convective surface Cahn–Hilliard equation and a generalized Navier-slip boundary condition. Compared to the models studied in [56, 32, 31], the Navier–Stokes–Cahn–Hilliard system introduced in [40] provides more regularity for the boundary quantities and therefore, the uniqueness of weak solutions can be established in two space dimensions. Moreover, due to the fourth-order dynamic boundary condition of Cahn–Hilliard type, the model in [40] is not only capable of overcoming the limitations (L1) and (L2) but also (L3).
In the present paper, we particularly want to consider the situation of creeping flows meaning that the Reynolds number
$\mathrm{Re}=\frac{uL}{\nu}$
associated with the fluids is very small ($\mathrm{Re}\ll 1$). This occurs if the flow speed $u$ and/or the characteristic length $L$ of the flow are small compared to the kinematic viscosity $\nu$. In this situation, it is not necessary to describe the time evolution of the velocity field by the full Navier–Stokes equation. Since advective inertial forces are small compared to viscous forces, the material derivative can be neglected. This leads to the Stokes equation. If a creeping flow through a porous medium is to be considered, an additional term accounting for the permeability needs to be included. The velocity field is then determined by the Brinkman equation.
Therefore, in this paper, we study the following Cahn–Hilliard–Brinkman system with dynamic boundary conditions:
$-\mathop{\rm div}\nolimits(2\nu(\varphi)\mathbf{D}{\boldsymbol{v}})+\lambda(\varphi){\boldsymbol{v}}+\nabla p=\mu\nabla\varphi\quad$ in $Q$, (1.1a)
$\mathop{\rm div}\nolimits({\boldsymbol{v}})=0\quad$ in $Q$, (1.1b)
$\partial_{t}\varphi+\mathop{\rm div}\nolimits(\varphi{\boldsymbol{v}})-\mathop{\rm div}\nolimits(M_{\Omega}(\varphi)\nabla\mu)=0\quad$ in $Q$, (1.1c)
$\mu=-\epsilon\Delta\varphi+\tfrac{1}{\epsilon}F^{\prime}(\varphi)\quad$ in $Q$, (1.1d)
$\partial_{t}\psi+\mathop{\rm div}\nolimits_{\Gamma}(\psi{\boldsymbol{v}})-\mathop{\rm div}\nolimits_{\Gamma}(M_{\Gamma}(\psi)\nabla_{\Gamma}\theta)=0\quad$ on $\Sigma$, (1.1e)
$\theta=-\epsilon_{\Gamma}\Delta_{\Gamma}\psi+\tfrac{1}{\epsilon_{\Gamma}}G^{\prime}(\psi)+\partial_{\bf n}\varphi\quad$ on $\Sigma$, (1.1f)
$K\partial_{\bf n}\varphi=\psi-\varphi\quad$ on $\Sigma$, (1.1g)
$M_{\Omega}(\varphi)\partial_{\bf n}\mu={\boldsymbol{v}}\cdot{\bf n}=0\quad$ on $\Sigma$, (1.1h)
$[2\nu(\varphi)\,\mathbf{D}{\boldsymbol{v}}\,{\bf n}+\gamma(\psi){\boldsymbol{v}}]_{\tau}=-[\psi\nabla_{\Gamma}\theta]_{\tau}\quad$ on $\Sigma$, (1.1i)
$\varphi(0)=\varphi_{0}\quad$ in $\Omega$, (1.1j)
$\psi(0)=\psi_{0}\quad$ on $\Gamma$. (1.1k)
[…]
In system (1), $\Omega\subset{\mathbb{R}}^{d}$ with $d\in\{2,3\}$ is a bounded domain with boundary $\Gamma:=\partial\Omega$, $T>0$ is a given final time, and for brevity, the notation $Q:=\Omega\times(0,T)$ and $\Sigma=\Gamma\times(0,T)$ is used. The vector-valued function ${\boldsymbol{v}}:Q\to{\mathbb{R}}^{d}$ stands for the volume averaged velocity field associated with the fluid mixture and
$\mathbf{D}{\boldsymbol{v}}=\frac{1}{2}\big{(}\nabla{\boldsymbol{v}}+(\nabla{\boldsymbol{v}})^{\top}\big{)}$
denotes the associated symmetric gradient. For the sake of simplicity, we will usually refrain from writing the trace operator. For instance, we will often simply write ${\boldsymbol{v}}$ instead of ${\boldsymbol{v}}|_{\Gamma}$. Nevertheless, in some instances, where confusion may arise, we will employ the explicit notation. For any vector field ${\boldsymbol{w}}$ on the boundary, we will write ${\boldsymbol{w}}_{\tau}:={\boldsymbol{w}}-({\boldsymbol{w}}\cdot{\bf n}){\bf n}$ to denote its tangential component. The symbols $\nabla_{\Gamma}$ and $\mathop{\rm div}\nolimits_{\Gamma}$ denote the surface gradient and the surface divergence, respectively, and $\Delta_{\Gamma}$ stands for the Laplace–Beltrami operator.
The functions $\varphi:Q\to{\mathbb{R}}$ and $\mu:Q\to{\mathbb{R}}$ denote the phase-field and the chemical potential in the bulk, respectively, whereas $\psi:\Sigma\to{\mathbb{R}}$ and $\theta:\Sigma\to{\mathbb{R}}$ represent the phase-field and the chemical potential on the boundary, respectively. Furthermore, the parameters $\epsilon$ and $\epsilon_{\Gamma}$ are positive real numbers which are related to the thickness of the diffuse interface in the bulk and on the surface, respectively. Therefore, these constants are usually chosen to be quite small. However, as their values do not have any impact on the mathematical analysis, we will simply fix $\epsilon=\epsilon_{\Gamma}=1$ in the subsequent sections. The phase-fields $\varphi$ and $\psi$ are directly related by the coupling condition (1.1g), where $K$ is a given nonnegative constant.
From a physical point of view, the kinematic viscosity $\nu(\varphi)$ and the permeability coefficient $\lambda(\varphi)$ in the Brinkman/Stokes equation (1.1a) can be expressed as
$\nu(\varphi)=\frac{\eta(\varphi)}{\varrho}\quad\text{and}\quad\lambda(\varphi)=\frac{\eta(\varphi)}{\varkappa},$
where $\eta(\varphi)>0$ denotes the dynamic viscosity, and the constants $\varrho>0$ and $\varkappa>0$ stand for the porosity and the intrinsic permeability of the porous medium, respectively. If both $\nu(\varphi)$ and $\lambda(\varphi)$ are positive, (1.1a) is the (quasi-stationary) Brinkman equation which describes the flow through a porous medium. However, if the porosity $\varkappa$ is large compared to the viscosity $\eta(\varphi)$, the function $\lambda(\varphi)$ is very small and can be neglected. In this case we enter the Stokes regime, where no porous media is considered (or the effects of the porous medium are at least negligible). In the formal limit $\varkappa\to\infty$ or $\lambda(\varphi)\to 0$, (1.1a) degenerates to the Stokes equation. In our analysis, we will be able to handle the Brinkman case ($\nu(\varphi)>0$ and $\lambda(\varphi)>0$) and the Stokes case ($\nu(\varphi)>0$ and $\lambda(\varphi)\equiv 0$) simultaneously. On the other hand, if $\lambda(\varphi)$ remains positive and the porosity $\varrho$ is large compared to the dynamic viscosity $\eta(\varphi)$ such that $\nu(\varphi)$ can be neglected, (1.1a) degenerates to Darcy’s law. However, we are not able to handle this case in terms of mathematical analysis as due to the absence of spatial derivatives of the velocity field in (1.1a), we would not obtain enough regularity to define the trace of ${\boldsymbol{v}}$ on the boundary in a reasonable manner.
The functions $F^{\prime}$ and $G^{\prime}$ are the derivatives of double-well potentials $F$ and $G$, respectively. Especially in applications related to materials science, a physically relevant choice for $F$ and/or $G$ is the logarithmic potential, which is also referred to as the Flory–Huggins potential. It is given by
$W_{\mathrm{log}}(s)=\frac{\Theta}{2}\big{[}(1+s)\,\ln(1+s)+(1-s)\,\ln(1-s)\big{]}+\frac{\Theta_{c}}{2}(1-s^{2}),$ (1.2)
[…]
$W_{\mathrm{pol}}(s)=\frac{\alpha}{4}(s^{2}-1)^{2}\quad\text{for all $s\in(-1,1)$},$ (1.3)
where $\alpha>0$ is a suitable constant. Another very commonly used singular potential is the double-obstacle potential, which is given by
$W_{\mathrm{obst}}(s)=\begin{cases}\frac{1}{2}(1-s^{2})&\text{if $|s|\leq 1$,}\\ +\infty&\text{else}.\end{cases}$ (1.4)
In the case $K=0$, the convective bulk-surface Cahn–Hilliard subsystem (1.1c)–(1.1h) is a special case of the one introduced in [40] since for the chemical potential $\mu$, a homogeneous Neumann type boundary condition is imposed in (1.1h). This corresponds to the choice $L=\infty$ in [40]. Therefore, by system (1), we describe a situation where no transfer of material between bulk and boundary occurs. However, its is important that due to the boundary conditions (1.1e)–(1.1i), the model (1) allows for dynamic changes of the contact angle as well as for a convection-induced motion of the contact line. This means that the limitations (L1) and (L2) explained above can be overcome. It is worth mentioning that this setup of dynamic boundary conditions for the Cahn–Hilliard equation (without coupling to a velocity equation) was originally derived in [50] by the Energetic Variational Approach. This system was further investigated in [19, 33, 53, 46]. For similar works on the Cahn–Hilliard equation with Cahn–Hilliard type dynamic boundary conditions, we refer to [49, 47, 34, 42, 14, 17, 18, 64].
In contrast to the model introduced in [40], the phase-fields $\varphi$ and $\psi$ are not just coupled by the trace relation $\varphi|_{\Sigma}=\psi$ on $\Sigma$, but by the more general Robin type coupling condition $K\partial_{\bf n}\varphi=\psi-\varphi$ with $K\geq 0$ (see (1.1g)). This also includes the trace relation via the choice $K=0$. The coupling condition (1.1g) was first used in [16] for an Allen–Cahn type dynamic boundary condition, and later in [46] for a Cahn–Hilliard type dynamic boundary condition. In particular, it was rigorously shown in [46] that the Dirichlet type coupling condition $\varphi|_{\Sigma}=\psi$ on $\Sigma$ can be recovered in the asymptotic limit $K\to 0$. From a physical point of view, the boundary condition (1.1g) with $K>0$ makes sense if the materials on the boundary may be different from those in the bulk. For instance, this might be the case if the materials on the boundary are transformed by chemical reactions. Apart from this, the boundary condition (1.1g) with $K>0$ has a crucial advantage for the mathematical analysis concerning the construction of weak solutions. The reason is that in the case $K=0$, the Dirichlet type coupling condition for the phase-fields already fixes one degree of freedom and is therefore a bad match for the no-mass-flux boundary condition (1.1h)1italic-($_{1}$. This seems to make the direct construction of weak solutions by a Faedo–Galerkin scheme impossible. However, employing (1.1g) with $K>0$, such problems do not arise. Hence, our strategy is to first prove the existence of weak solutions for $K>0$ by a Faedo–Galerkin approach. Afterwards, we construct a weak solution associated with $K=0$ by passing to the limit $K\to 0$ in a suitable sense.
An important property of the system (1) (for any $K\geq 0$) is its thermodynamic consistency with respect to the free energy functional
$E_{K}(\varphi,\psi)$ $:=\int_{\Omega}\Big{(}\frac{\epsilon}{2}|\nabla\varphi|^{2}+\frac{1}{\epsilon}F(\varphi)\Big{)}+\int_{\Gamma}\Big{(}\frac{\epsilon_{\Gamma}}{2}|\nabla_{\Gamma}\psi|^{2}+\frac{1}{\epsilon_{\Gamma}}G(\psi)\Big{)}$
$\quad+\frac{\sigma(K)}{2}\int_{\Gamma}(\psi-\varphi)^{2},$ (1.5)
where $\sigma(K)=K^{-1}$ if $K>0$ and $\sigma(K)=0$ if $K=0$. This means that sufficiently regular solutions of (1) satisfy the energy dissipation law
[…]
on $[0,T]$, where all the terms on the right-hand side are non-positive and can be interpreted as the dissipation rate. Compared to the model in [40], the additional term $\int_{\Omega}\lambda(\varphi)\left|{\boldsymbol{v}}\right|^{2}$ arises due to dissipative effects caused by the porous medium.
As mentioned above, due to the usage of the no-mass-flux condition $M_{\Omega}(\varphi)\partial_{\bf n}\mu=0$ on $\Sigma$ (see (1.1h)), we do not describe any transfer of material between bulk and surface. This entails that the bulk mass and the surface mass are conserved separately, i.e., sufficiently regular solutions satisfy the mass conservation laws
$\frac{1}{|\Omega|}\int_{\Omega}\varphi(t)$ $=\frac{1}{|\Omega|}\int_{\Omega}\varphi_{0}=:m_{0}$ (1.6)
$\frac{1}{|\Gamma|}\int_{\Gamma}\psi(t)$ $=\frac{1}{|\Gamma|}\int_{\Gamma}\psi_{0}=:m_{\Gamma 0}$ (1.7)
for all $t\in[0,T]$.
The structure of the paper is as follows. In Section 2 we collect some notation, assumptions, preliminaries and important tools as well as the main results of our work. In the case of regular potentials, the Cahn–Hilliard–Brinkman system is analyzed in Section 3. There, we first establish the existence of a weak solution in the case $K>0$, where phase-fields are coupled by a Robin type boundary condition. By sending $K\to 0$, we then obtain the existence of a weak solution also in the case $K=0$, where the phase-fields are coupled by a trace relation. Afterwards, we prove uniqueness in all cases $K\geq 0$ under suitable additional assumptions. Eventually, in Section 4, we show the existence of a weak solution to the Cahn–Hilliard–Brinkman system with singular potentials. This is done based on the results established in Section 3 by approximating the singular potentials by means of Yosida regularizations and finally passing to the limit.
2   Preliminaries and main results
2.1   Notation
Throughout the manuscript, $\Omega$ is a bounded domain in ${\mathbb{R}}^{d}$, $d\in\{2,3\}$, with Lipschitz boundary $\Gamma:=\partial\Omega$ and ${\bf n}$ is the associated outward unit normal vector field. We write $|\Omega|$ and $|\Gamma|$ to denote the Lebesgue measure of $\Omega$ and the Hausdorff measure of $\Gamma$, respectively. For any given Banach space $X$, we denote its norm by $\left\|\,\cdot\,\right\|_{X}$, its dual space by $X^{*}$ and the duality pairing between $X^{*}$ and $X$ by $\mathopen{\langle}\cdot,\cdot\mathclose{\rangle}_{X}$. Besides, if $X$ is a Hilbert space, we write $(\cdot,\cdot)_{X}$ to denote the corresponding inner product. For every $1\leq p\leq\infty$, $k\geq 0$ and $s>0$, the standard Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces defined on $\Omega$ are denoted by $L^{p}(\Omega)$, $W^{k,p}(\Omega)$ and $H^{s}(\Omega)$, and their standard norms are denoted by $\left\|\,\cdot\,\right\|_{L^{p}(\Omega)}$ $\left\|\,\cdot\,\right\|_{W^{k,p}(\Omega)}$ and $\left\|\,\cdot\,\right\|_{H^{s}(\Omega)}$, respectively. It is well known that the spaces $H^{0}(\Omega)=L^{2}(\Omega)$ and $H^{k}(\Omega)=W^{k,2}(\Omega)$ for all $k\in{\mathbb{N}}$ can be identified, and these spaces are Hilbert spaces. The Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces on the boundary $\Gamma$ are defined analogously. For more details, we refer to [62, 63]. Moreover, for any Banach spaces $X$ and $Y$, their intersection $X\cap Y$ is also a Banach space subject to the norm
$\left\|v\right\|_{X\cap Y}:=\left\|v\right\|_{X}+\left\|v\right\|_{Y},\quad v\in X\cap Y.$
As some spaces will appear very frequently, we introduce the shortcuts:
$H$ $:=L^{2}(\Omega),\quad$ $H_{\Gamma}$ $:=L^{2}(\Gamma),\quad$ $V$ $:=H^{1}(\Omega),\quad$ ${V}_{\Gamma}$ $:=H^{1}(\Gamma),$
${\bf{H}}$ $:=L^{2}(\Omega;{\mathbb{R}}^{d}),\quad$ $\mathbf{H}_{\Gamma}$ $:=L^{2}(\Gamma;{\mathbb{R}}^{d}),\quad$ ${\bf{V}}$ $:=H^{1}(\Omega;{\mathbb{R}}^{d}),$
${\mathbb{H}}$ $:=L^{2}(\Omega;{\mathbb{R}}^{d\times d}).$
We further introduce the spaces of solenoidal (divergence-free) velocity fields:
${\bf H}_{\sigma,{\bf n}}$ :=$:=\big{\{}{\boldsymbol{w}}\in{\bf{H}}:\mathop{\rm div}\nolimits({\boldsymbol{w}})=0\;\text{in $\Omega$},\;{\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}=0\;\text{on $\Gamma$}\big{\}},$
${\bf V}_{\sigma,{\bf n}}$ $:={\bf{V}}\cap{\bf H}_{\sigma,{\bf n}}.$
We point out that in the definition of ${\bf H}_{\sigma,{\bf n}}$, the relation $\mathop{\rm div}\nolimits({\boldsymbol{w}})=0$ in $\Omega$ is to be understood in the sense of distributions. This already implies ${\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}\in H^{-1/2}(\Gamma)$, and therefore, the relation ${\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}=0$ on $\Gamma$ is well-defined. As ${\bf H}_{\sigma,{\bf n}}$ and ${\bf V}_{\sigma,{\bf n}}$ are closed linear subspaces of the Hilbert spaces ${\bf{H}}$ and ${\bf{V}}$, respectively, they are also Hilbert spaces. We further introduce the bulk-surface product spaces
${\cal H}$ $:=H\times H_{\Gamma},\quad{\cal V}:=V\times{V}_{\Gamma},$
${\cal V}_{K}$ :=$:=\begin{cases}{\cal V}&\text{if $K>0$},\\ \{(w,w_{\Gamma})\in{\cal V}:w_{\Gamma}={w}|_{\Gamma}\,\,\text{on $\Gamma$}\}&\text{if $K=0$},\end{cases}$
and endow them with the corresponding inner products
$\big{(}(v,v_{\Gamma}),(w,w_{\Gamma})\big{)}_{\cal H}$ $:=(v,w)_{H}+{(v_{\Gamma},w_{\Gamma})}_{H_{\Gamma}}$ for all $(v,v_{\Gamma}),(w,w_{\Gamma})\in{{\cal H}}$,
$\big{(}(v,v_{\Gamma}),(w,w_{\Gamma})\big{)}_{\cal V}$ $:=(v,w)_{V}+{(v_{\Gamma},w_{\Gamma})}_{{V}_{\Gamma}}$ for all $(v,v_{\Gamma}),(w,w_{\Gamma})\in{\cal V}$,
so that ${\cal H}$, ${\cal V}$ and ${\cal V}_{K}$ are Hilbert spaces. For any $v\in{V^{*}}$ and $v_{\Gamma}\in{V}_{\Gamma}^{*}$, we define the generalized mean values by
$\mathopen{\langle}v\mathclose{\rangle}_{\Omega}:=\frac{1}{|\Omega|}\,\mathopen{\langle}v,1\mathclose{\rangle}_{V}\,,\quad\mathopen{\langle}v_{\Gamma}\mathclose{\rangle}_{\Gamma}:=\frac{1}{|\Gamma|}\,\mathopen{\langle}v_{\Gamma},1\mathclose{\rangle}_{{V}_{\Gamma}}\,,$ (2.1)
where $1$ represents the constant function assuming value $1$ in $\Omega$ and on $\Gamma$, respectively. To introduce a weak formulation of (1), it will be useful to define the function
$\sigma:[0,\infty)\to[0,\infty),\quad\sigma(r)=\begin{cases}\frac{1}{r}&\text{if $r>0$},\\ 0&\text{if $r=0$}\end{cases}$ (2.2)
to handle the cases $K>0$ and $K=0$ simultaneously.
2.2   General assumptions
  1. (A1)
    The set $\Omega\subset{\mathbb{R}}^{d}$ with $d\in\{2,3\}$ is a bounded Lipschitz domain.
  2. (A2)
    The mobility functions $M_{\Omega}:{\mathbb{R}}\to{\mathbb{R}}$ and $M_{\Gamma}:{\mathbb{R}}\to{\mathbb{R}}$ are continuous, bounded and uniformly positive. This means that there exist positive constants $M_{1}$, $M_{2}$, $M_{\Gamma,1}$ and $M_{\Gamma,2}$ such that
    $0<M_{1}\leq M_{\Omega}(r)\leq M_{2},\quad 0<M_{\Gamma,1}\leq M_{\Gamma}(r)\leq M_{\Gamma,2}\quad\text{for all $r\in{\mathbb{R}}.$}$
  3. (A3)
    The viscosity function $\nu:{\mathbb{R}}\to{\mathbb{R}}$ and the friction parameter $\gamma:{\mathbb{R}}\to{\mathbb{R}}$ are continuous, bounded and uniformly positive. Namely, there exist positive constants $\nu_{1}$, $\nu_{2}$, $\gamma_{1}$ and $\gamma_{2}$ such that
    $0<\nu_{1}\leq\nu(r)\leq\nu_{2},\quad 0<\gamma_{1}\leq\gamma(r)\leq\gamma_{2}\quad\text{for all $r\in{\mathbb{R}}.$}$
    Furthermore, the permeability function $\lambda:{\mathbb{R}}\to{\mathbb{R}}$ is continuous, bounded and nonnegative, that is, there exists a nonnegative constant $\lambda_{1}$ such that
    $0\leq\lambda(r)\leq\lambda_{1}\quad\text{for all $r\in{\mathbb{R}}.$}$
2.3   Preliminaries
Throughout the paper, we will frequently use the following bulk-surface Poincaré inequality:
Lemma 2.1.
There exists a constant $C_{P}^{\prime}$ depending only on $\Omega$ such that
$\|u\|_{H}\leq C_{P}^{\prime}\left(\|\nabla u\|_{{\bf H}}+\|u\|_{{H}_{\Gamma}}\right)\quad\text{for all $u\in V$.}$ (2.3)
A proof of this inequality can be found, for instance, in [61, Chapter II, Section 1.4].
We further recall the following result on interpolation between Sobolev spaces:
Lemma 2.2.
Let $U\subset{\mathbb{R}}^{m}$ with $m\in\mathbb{N}$ be a bounded Lipschitz domain, and suppose that $\theta\in(0,1)$ and $r,s_{0},s_{1}\in{\mathbb{R}}$ satisfy
$r=(1-\theta)s_{0}+\theta s_{1}.$
We further assume that $U$ is of class $C^{\ell}$ with an integer $\ell\geq\max\{s_{0},s_{1}\}$. Then, there exist positive constants $C_{U}$ and $C_{\partial U}$ depending only on $U$, $r$, $s_{0}$, $s_{1}$ and $\theta$ such that the following interpolation inequalities hold:
$\|f\|_{H^{r}(U)}$ $\leq C_{U}\|f\|_{H^{s_{0}}(U)}^{1-\theta}\|f\|_{H^{s_{1}}(U)}^{\theta},$ (2.4)
$\|f\|_{H^{r}(\partial U)}$ $\leq C_{\partial U}\|f\|_{H^{s_{0}}(\partial U)}^{1-\theta}\|f\|_{H^{s_{1}}(\partial U)}^{\theta}.$ (2.5)
Inequality (2.4) follows from an interpolation result shown in [62, Sec 4.3.1, Theorem 1 and Remark 1], whereas (2.5) follows from an interpolation result presented in [63, Sec 7.4.5, Remark 2].
2.4   The Cahn–Hilliard–Brinkman system with regular potentials
First, we present our mathematical results for system (1) in the case of regular double-well potentials $F$ and $G$. As mentioned above, we simply set $\epsilon=\epsilon_{\Gamma}=1$, as the exact values of these interface parameters do not have any impact on the mathematical analysis as long as they are positive.
2.4.1   Assumptions for regular potentials
  1. (R1)
    The potentials $F:{\mathbb{R}}\to[0,\infty)$ and $G:{\mathbb{R}}\to[0,\infty)$ are continuously differentiable, and there exist exponents $p,q\in{\mathbb{R}}$ with
    $p\in\begin{cases}[2,\infty)&\text{if $d=2$,}\\ [2,6]&\text{if $d=3$,}\end{cases}\quad\text{and}\quad q\in[2,\infty)$ (2.6)
    as well as constants $c_{F^{\prime}},c_{G^{\prime}}\geq 0$ such that
    $\left|F^{\prime}(r)\right|$ $\leq c_{F^{\prime}}(1+\left|r\right|^{p-1}),$ (2.7)
    $\left|G^{\prime}(r)\right|$ $\leq c_{G^{\prime}}(1+\left|r\right|^{q-1})$ (2.8)
    for all $r\in{\mathbb{R}}$. This implies that there exist constants $c_{F},c_{G}\geq 0$ such that $F$ and $G$ fulfill the growth conditions
    $F(r)$ $\leq c_{F}(1+\left|r\right|^{p}),$ (2.9)
    $\quad G(r)$ $\leq c_{G}(1+\left|r\right|^{q})$ (2.10)
    for all $r\in{\mathbb{R}}$.
  2. (R2)
    In addition to (R1), $F$ and $G$ are twice continuously differentiable and there exist constants $c_{F^{\prime\prime}},c_{G^{\prime\prime}}\geq 0$ such that
    $\left|F^{\prime\prime}(r)\right|$ $\leq c_{F^{\prime\prime}}(1+\left|r\right|^{p-2}),$ (2.11)
    $\left|G^{\prime\prime}(r)\right|$ $\leq c_{G^{\prime\prime}}(1+\left|r\right|^{q-2})$ (2.12)
    for all $r\in{\mathbb{R}}$, where $p$ and $q$ are the exponents introduced in (2.6).
2.4.2   Definition of weak solutions for regular potentials
Definition 2.3.
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. A quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1) if the following conditions are fulfilled:
  1. (i)
    The functions ${\boldsymbol{v}}$, $\varphi$, $\mu$, $\psi$ and $\theta$ have the regularity
    ${\boldsymbol{v}}$ $\in L^{2}(0,T;{\bf V}_{\sigma,{\bf n}}),\quad{\boldsymbol{v}}|_{\Gamma}\in L^{2}(0,T;{\bf H}_{\Gamma}),$
    $(\varphi,\psi)$ $\in H^{1}(0,T;{{\cal V}_{K}^{*}})\cap C^{0}([0,T];{\cal H})\cap L^{\infty}(0,T;{\cal V}_{K}),$
    $(\mu,\theta)$ $\in L^{2}(0,T;{\cal V}).$
  2. (ii)
    The variational formulation
[…]
    $\mathopen{\langle}\partial_{t}\varphi,\zeta\mathclose{\rangle}_{V}-\int_{\Omega}\varphi{\boldsymbol{v}}\cdot\nabla\zeta+\int_{\Omega}M_{\Omega}(\varphi)\nabla\mu\cdot\nabla\zeta=0,$ (2.13b)
    $\mathopen{\langle}\partial_{t}\psi,{\zeta_{\Gamma}}\mathclose{\rangle}_{{V}_{\Gamma}}-\int_{\Gamma}\psi{\boldsymbol{v}}\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}+\int_{\Gamma}M_{\Gamma}(\psi)\nabla_{\Gamma}\theta\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}=0,$ (2.13c)
[…]
    holds a.e. in $[0,T]$ for all ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, ${\zeta_{\Gamma}}\in V_{\Gamma}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}_{K}$.
  3. (iii)
    The initial conditions are satisfied in the follwing sense:
[…]
  4. (iv)
    The weak energy dissipation law
    $E_{K}\big{(}\varphi(t),\psi(t)\big{)}+2\int_{0}^{t}\int_{\Omega}\nu(\varphi)|\mathbf{D}{\boldsymbol{v}}|^{2}+\int_{0}^{t}\int_{\Omega}\lambda(\varphi)|{\boldsymbol{v}}|^{2}+\int_{0}^{t}\int_{\Gamma}\gamma(\psi)|{\boldsymbol{v}}|^{2}$
    $\qquad+\int_{0}^{t}\int_{\Omega}M_{\Omega}(\varphi)|\nabla\mu|^{2}+\int_{0}^{t}\int_{\Gamma}M_{\Gamma}(\psi)|\nabla_{\Gamma}\theta|^{2}$
    $\quad\leq E_{K}(\varphi_{0},\psi_{0})$ (2.14)
    holds for all $t\in[0,T]$.
2.4.3   Existence of a weak solution in the case $K>0$
We first show the existence of a weak solution to the Cahn–Hilliard–Brinkman system (1) in the case $K>0$.
Theorem 2.4.
Let $K>0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. Then, the Cahn–Hilliard–Brinkman system (1) possesses at least one weak solution $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ in the sense of Definition 2.3, which further satisfies $(\mu,\theta)\in L^{4}(0,T;{\cal H})$.
Let us now assume that the domain $\Omega$ is of class $C^{\ell}$ with $\ell\in\{2,3\}$. If $d=3$, we further assume $p<6$, and if $\ell=3$, we further assume that (R2) holds. Then, we have the additional regularities
$(\varphi,\psi)$ $\in L^{4}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)}\qquad\text{in case $\ell\in\{2,3\}$},$ (2.15a)
$(\varphi,\psi)$ $\in L^{2}\big{(}0,T;H^{3}(\Omega)\times H^{3}(\Gamma)\big{)}\qquad\text{in case $\ell=3$},$ (2.15b)
and the equations (1.1d), (1.1f) and (1.1g) are fulfilled in the strong sense, that is, almost everywhere in $Q$ and on $\Sigma$, respectively. In the case $\ell=3$, we further have
$(\varphi,\psi)\in C^{0}\big{(}[0,T];{\cal V}\big{)}.$ (2.16)
2.4.4   The limit $K\to 0$ and existence of a weak solution in the case $K=0$
We now investigate the limit $K\to 0$ in which the boundary condition (1.1g) formally tends to the Dirichlet condition $\psi=\varphi|_{\Gamma}$ almost everywhere on $\Sigma$. In the following theorem, we send $K\to 0$ in system (1) to prove the existence of a weak solution to (1) in the case $K=0$, and we further specify the convergence properties of this asymptotic limit.
Theorem 2.5.
Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{0}$ be any initial data. Let $({K_{n}})_{n\in{\mathbb{N}}}$ be a sequence of positive real numbers such that ${K_{n}}\to 0$ as $n\to\infty$. For any $n\in{\mathbb{N}}$, let $({\boldsymbol{v}}^{K_{n}},\varphi^{K_{n}},\mu^{K_{n}},\psi^{K_{n}},\theta^{K_{n}})$ denote any weak solution corresponding to ${K_{n}}>0$ in the sense of Definition 2.3. Then, there exists a quintuplet of functions $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ with $\varphi^{0}|_{\Gamma}=\psi^{0}$ a.e. on $\Sigma$ such that for any $s\in[0,1)$,
${\boldsymbol{v}}^{K_{n}}$ $\to{\boldsymbol{v}}^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (2.17a)
${\boldsymbol{v}}^{K_{n}}|_{\Gamma}$ $\to{\boldsymbol{v}}^{0}|_{\Gamma}$ weakly in $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (2.17b)
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$ (2.17c)
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$ (2.17d)
$\mu^{K_{n}}$ $\to\mu^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;V)$},$ (2.17e)
$\theta^{K_{n}}$ $\to\theta^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})$},$ (2.17f)
$\varphi^{K_{n}}|_{\Gamma}-\psi^{K_{n}}$ $\to 0$ strongly in L(0,T;HΓ), and a.e. on $\quad\text{strongly in $L^{\infty}(0,T;{H}_{\Gamma})$, and {a.e.\ on~{}$\Sigma$}},$ (2.17g)
as $n\to\infty$ along a non-relabeled subsequence. Moreover, the limit $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman model (1) in the sense of Definition 2.3 with $K=0$.
Let us now assume that the domain $\Omega$ is of class $C^{\ell}$ with $\ell\in\{2,3\}$. If $d=3$, we further assume $p\leq 4$, and if $\ell=3$, we further assume that (R2) holds. Then, we have the additional regularities
$(\varphi^{0},\psi^{0})$ $\in L^{2}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)}\qquad\text{in case $\ell\in\{2,3\}$},$ (2.18a)
$(\varphi^{0},\psi^{0})$ $\in L^{2}\big{(}0,T;H^{3}(\Omega)\times H^{3}(\Gamma)\big{)}\qquad\text{in case $\ell=3$},$ (2.18b)
and the equations (1.1d) and (1.1f) are fulfilled in the strong sense. Moreover, in the case $\ell=3$, we further have
$(\varphi^{0},\psi^{0})$ $\in C^{0}\big{(}[0,T];{\cal V}_{0}\big{)}\cap L^{4}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)},$ (2.19)
$(\mu^{0},\theta^{0})$ $\in L^{4}\big{(}0,T;{\cal H}\big{)}.$ (2.20)
Remark 2.6.
The additional assumption $p\leq 4$ in the second paragraph of the above theorem was merely imposed to avoid unnecessary technicalities in the proof for additional regularity of the phase-fields. However, it should be possible to obtain the same regularities also for $p\in(4,6)$. This could be established by approximating the potential $F$ by a sequence $(F_{k})_{k\in{\mathbb{N}}}$ of potentials satisfying (R1) with $p\leq 4$. Then, proceeding similarly as in the proof of Theorem 2.4 (see Subsection 3.1.5) additional uniform bounds on the corresponding weak solutions $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ would have to be derived to obtain the weak convergences $(\varphi_{k},\psi_{k})\to(\varphi,\psi)$ in $L^{2}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)}$ in case $\ell\in\{2,3\}$, and $(\varphi_{k},\psi_{k})\to(\varphi,\psi)$ in $L^{2}\big{(}0,T;H^{3}(\Omega)\times H^{3}(\Gamma)\big{)}$ in case $\ell=3$, which imply the desired regularities.
2.4.5   Stability and uniqueness of the weak solution in the general case $K\geq 0$
In the case of regular potentials, constant mobilities, and a constant viscosity, we are able to prove the uniqueness of the weak solutions established in Theorem 2.4 provided that the following assumption on the potentials $F$ and $G$ holds.
  1. (Renumi)
    The potentials $F:{\mathbb{R}}\to[0,\infty)$ and $G:{\mathbb{R}}\to[0,\infty)$ are three times continuously differentiable, and there exist exponents $p,q\in{\mathbb{R}}$ with
    $p\in\begin{cases}[3,\infty)&\text{if $d=2$,}\\ [3,6)&\text{if $d=3$ and $K>0$,}\\ [3,4]&\text{if $d=3$ and $K=0$,}\end{cases}\qquad\text{and}\qquad q\in[3,\infty)$ (2.21)
    as well as constants $c_{F^{(3)}},c_{G^{(3)}}\geq 0$ such that the third-order derivatives satisfy
    $\left|F^{(3)}(r)\right|$ $\leq c_{F^{(3)}}(1+\left|r\right|^{p-3}),$ (2.22)
    $\left|G^{(3)}(r)\right|$ $\leq c_{G^{(3)}}(1+\left|r\right|^{q-3})$ (2.23)
    for all $r\in{\mathbb{R}}$.
We point out that (Renumi) implies (R1) and (R2) with $p$ and $q$ being chosen as in (2.21). The restriction $p\in[3,4]$ in the case $d=3$ and $K=0$ is imposed since in Theorem 2.5, higher regularity properties for the phase-fields were established only for $p\leq 4$. However, as pointed out in Remark 2.6, it should be possible to relieve this restriction such that also in the case $d=3$ and $K=0$, merely $p\in[3,6)$ would have to be imposed in (Renumi).
Theorem 2.7.
Suppose that (A1) and (Renumi) are fulfilled and let $K\geq 0$ be arbitrary. In addition to (A2) and (A3), we further assume that $\gamma$ and $\lambda$ are Lipschitz continuous functions and that the functions $\nu$, $M_{\Omega}$ and $M_{\Gamma}$ reduce to positive constants denoted by the same symbols. For any $i\in\{1,2\}$, let $(\varphi_{0,i},\psi_{0,i})\in{\cal V}_{K}$ be any pair of initial data, and let $({\boldsymbol{v}}_{i},\varphi_{i},\mu_{i},\psi_{i},\theta_{i})$ be a corresponding weak solution in the sense of Definition 2.3. Then, the stability estimate
$\left\|{\boldsymbol{v}}_{1}-{\boldsymbol{v}}_{2}\right\|_{L^{2}(0,T;{\bf{V}})}+\left\|\varphi_{1}-\varphi_{2}\right\|_{L^{\infty}(0,T;V)}+\left\|\mu_{1}-\mu_{2}\right\|_{L^{2}(0,T;V)}+\left\|\psi_{1}-\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}$
$\quad+\left\|\theta_{1}-\theta_{2}\right\|_{L^{2}(0,T;{V}_{\Gamma})}\leq C_{S}\big{(}\left\|\varphi_{0,1}-\varphi_{0,2}\right\|_{V}+\left\|\psi_{0,1}-\psi_{0,2}\right\|_{{V}_{\Gamma}}\big{)}$ (2.24)
holds for a constant $C_{S}\geq 0$ depending only on $K$, $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3) and (Renumi). In particular, choosing $(\varphi_{0,1},\psi_{0,1})=(\varphi_{0,2},\psi_{0,2})$, this entails the uniqueness of the corresponding weak solution.
2.5   The Cahn–Hilliard–Brinkman system with singular potentials
We now consider the system (1) for a general class of singular potentials. For those, we manage to establish just existence of weak solutions due to the lower regularity at disposal. Recall that $\epsilon=\epsilon_{\Gamma}=1$ as mentioned above.
2.5.1   Assumptions for singular potentials
For the potentials $F$ and $G$, we now make the following assumptions.
  1. (S1)
    The potentials $F$ and $G$ can be decomposed as $F=\widehat{\beta}+\widehat{\pi}$ and $G={\widehat{\beta}}_{\Gamma}+{\widehat{\pi}}_{\Gamma}$.
    Here, $\widehat{\beta},{\widehat{\beta}}_{\Gamma}:{\mathbb{R}}\to[0,\infty]$ are lower semicontinuous and convex functions with $\widehat{\beta}(0)=0$ and ${\widehat{\beta}}_{\Gamma}(0)=0$. For brevity, we define
    $\beta:=\partial\widehat{\beta}\quad\text{and}\quad\beta_{\Gamma}:=\partial{\widehat{\beta}_{\Gamma}}.$
    Moreover, we suppose that $\widehat{\pi},{\widehat{\pi}}_{\Gamma}\in C^{1}({\mathbb{R}})$ with Lipschitz continuous derivatives $\pi:=\widehat{\pi}^{\prime}$ and $\pi_{\Gamma}:={\widehat{\pi}}_{\Gamma}^{\prime}$.
    We point out that $\beta$ and ${\beta_{\Gamma}}$ are maximal monotone graphs in ${\mathbb{R}}\times{\mathbb{R}}$ whose effective domains are denoted by $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. In particular, as 00 is a minimum point of both $\widehat{\beta}$ and ${\widehat{\beta}}_{\Gamma}$, it turns out that $0\in\beta(0)$ and $0\in{\beta_{\Gamma}}(0)$. Finally, we denote by $\beta^{\circ}$ the minimal section of the graph $\beta$, which is defined as
    $\beta^{\circ}(r)\coloneqq\Big{\{}r^{\ast}\in D(\beta):\left|r^{\ast}\right|=\min_{s\in\beta(r)}\left|s\right|\Big{\}}\qquad\text{for all $r\in D(\beta)$}$
    (see, e.g., [12]). The same definition applies to $\beta^{\circ}_{\Gamma}$ for ${\beta_{\Gamma}}$.
  2. (S2)
    We also assume the growth condition
    $\lim_{r\to+\infty}\frac{\widehat{\beta}(r)}{|r|^{2}}=+\infty\,.$ (2.25)
    Moreover, we demand $D(\beta_{\Gamma})\subseteq D(\beta)$, and postulate that the boundary graph dominates the bulk graph in the following sense:
    $\exists\,\kappa_{1},\kappa_{2}>0:\quad|\beta^{\circ}(r)|\leq\kappa_{1}|\beta_{\Gamma}^{\circ}(r)|+\kappa_{2}\quad\text{for every $r\in D(\beta_{\Gamma})$.}$ (2.26)
    Here, $\beta^{\circ}$ and $\beta_{\Gamma}^{\circ}$ are the minimal sections introduced in (S1).
Note that all the examples of potentials given in (1.2)–(1.4) fulfill the assumptions (S1) and (S2), provided that the boundary potential dominates the one in the bulk as demanded in (2.26). In particular, the only scenario where a singular and a regular potential may coexist is the case in which the boundary potential is the singular one. This assumption has first been made in [13] and was used afterwards in several contributions in the literature, see, e.g., [15, 21, 24, 23, 25, 22, 17, 18, 19, 14]. However, in some other works such as [36, 37] different compatibility conditions were assumed.
2.5.2   Definition of weak solutions for singular potentials
Definition 2.8.
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3), (S1) and (S2) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data satisfying
$\widehat{\beta}(\varphi_{0})$ $\in L^{1}(\Omega),\quad$ $m_{0}:=\mathopen{\langle}\varphi_{0}\mathclose{\rangle}_{\Omega}\in{\rm int}(D(\beta)),$ (2.27a)
$\widehat{\beta}_{\Gamma}(\psi_{0})$ $\in L^{1}(\Gamma),\quad$ $m_{\Gamma 0}:=\mathopen{\langle}\psi_{0}\mathclose{\rangle}_{\Gamma}\in{\rm int}(D(\beta_{\Gamma})).$ (2.27b)
Then, $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1) if the following conditions are fulfilled:
  1. (i)
    The functions ${\boldsymbol{v}}$, $\varphi$, $\xi$, $\mu$, $\psi$, $\xi_{\Gamma}$ and $\theta$ have the regularity
    ${\boldsymbol{v}}$ $\in L^{2}(0,T;{\bf V}_{\sigma,{\bf n}}),\quad{\boldsymbol{v}}|_{\Gamma}\in L^{2}(0,T;{\bf H}_{\Gamma}),$
    $(\varphi,\psi)$ $\in H^{1}(0,T;{\cal V}_{K}^{*})\cap C^{0}([0,T];{{\cal H}})\cap L^{\infty}(0,T;{\cal V}_{K}),\quad$
    $(\xi,\xi_{\Gamma})$ $\in L^{2}(0,T;{\cal H}),$
    $(\mu,\theta)$ $\in L^{2}(0,T;{\cal V}).$
  2. (ii)
    The variational formulation
[…]
    $\mathopen{\langle}\partial_{t}\varphi,\zeta\mathclose{\rangle}_{V}-\int_{\Omega}\varphi{\boldsymbol{v}}\cdot\nabla\zeta+\int_{\Omega}M_{\Omega}(\varphi)\nabla\mu\cdot\nabla\zeta=0,$ (2.28b)
    $\mathopen{\langle}\partial_{t}\psi,{\zeta_{\Gamma}}\mathclose{\rangle}_{{V}_{\Gamma}}-\int_{\Gamma}\psi{\boldsymbol{v}}\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}+\int_{\Gamma}M_{\Gamma}(\psi)\nabla_{\Gamma}\theta\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}=0,$ (2.28c)
[…]
    holds a.e. in $[0,T]$ for all ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, ${\zeta_{\Gamma}}\in V_{\Gamma}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}_{K}$, where
[…]
  3. (iii)
    The initial conditions are satisfied in the follwing sense:
[…]
  4. (iv)
    The weak energy dissipation law
    $E_{K}\big{(}\varphi(t),\psi(t)\big{)}+2\int_{0}^{t}\int_{\Omega}\nu(\varphi)|\mathbf{D}{\boldsymbol{v}}|^{2}+\int_{0}^{t}\int_{\Omega}\lambda(\varphi)|{\boldsymbol{v}}|^{2}+\int_{0}^{t}\int_{\Gamma}\gamma(\psi)|{\boldsymbol{v}}|^{2}$
    $\qquad+\int_{0}^{t}\int_{\Omega}M_{\Omega}(\varphi)|\nabla\mu|^{2}+\int_{0}^{t}\int_{\Gamma}M_{\Gamma}(\psi)|\nabla_{\Gamma}\theta|^{2}$
    $\quad\leq E_{K}(\varphi_{0},\psi_{0})$
    holds for all $t\in[0,T]$.
2.5.3   Existence of a weak solution
Theorem 2.9.
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (S1)(S2) are fulfilled. Let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ denote any initial data satisfying (2.8). In the case $K=0$, let the domain $\Omega$ be of class $C^{2}$. Then, the Cahn–Hilliard–Brinkman system (1) admits at least one weak solution $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ in the sense of Definition 2.8. In all cases, if the domain $\Omega$ is at least of class $C^{2}$, it holds that
$\varphi\in L^{2}(0,T;H^{2}(\Omega)),\quad\psi\in L^{2}(0,T;H^{2}(\Gamma))$ (2.29)
and the equations
$\mu$ $=-\Delta\varphi+\xi+\pi(\varphi)\quad$ in $Q$, (2.30)
$\theta$ $=-\Delta_{\Gamma}\psi+\xi_{\Gamma}+\pi_{\Gamma}(\psi)+\partial_{\bf n}\varphi\quad$ on $\text{on $\Sigma$},$ (2.31)
$K\partial_{\bf n}\varphi$ $=\varphi-\psi\quad$ on $\Sigma$ (2.32)
are fulfilled in the strong sense.
3   Analysis of the Cahn–Hilliard–Brinkman system with regular potentials
3.1   Existence of weak solutions in the case $K>0$
Proof of Theorem 2.4.
We intend to construct a weak solution to system (1) by discretizing the weak formulation ((ii)) by means of a Faedo–Galerkin scheme. In this proof, the letter $C$ will denote generic positive constants that may depend on $K$, $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3), and may change their value from line to line. Recall that, as $K$ is assumed to be positive here, we have $\sigma(K)=\frac{1}{K}$.
3.1.1   Construction of local-in-time approximate solutions
It is well known that the Poisson–Neumann eigenvalue problem
$-\Delta u=\lambda_{\Omega}u\;\;\text{in $\Omega$},\quad\partial_{\bf n}u=0\;\;\text{on $\Gamma$}$ (3.1)
possesses countably many eigenvalues and a corresponding sequence $\{u_{i}\}_{i\in\mathbb{N}}\subset V$ of $H$-normalized eigenfunctions which form an orthonormal basis of $H$ and an orthogonal Schauder basis of $V$. Similarly, invoking the spectral theorem for compact self-adjoint operators, it follows that the eigenvalue problem
$-\Delta_{\Gamma}v=\lambda_{\Gamma}v\;\;\text{on $\Gamma$}$ (3.2)
for the Laplace–Beltrami operator possesses countably many eigenvalues and a corresponding sequence $\{v_{i}\}_{i\in\mathbb{N}}\subset V_{\Gamma}$ of $H_{\Gamma}$-normalized eigenfunctions which form an orthonormal basis of $H_{\Gamma}$ and an orthogonal Schauder basis of $V_{\Gamma}$. For any $k\in\mathbb{N}$, we now define the finite-dimensional subspaces
$V_{k}$ $:={\rm span}\{u_{i}:1\leq i\leq k\}\subset V,$
$V_{\Gamma,k}$ $:={\rm span}\{v_{j}:1\leq j\leq k\}\subset V_{\Gamma},$
${\cal V}_{k}$ $:={\rm span}\{(u_{i},v_{j}):1\leq i,j\leq k\}\subset{\cal V}.$
We point out that, due to the above considerations, the inclusions
$\bigcup_{k\in\mathbb{N}}V_{k}\subseteq V,\qquad\bigcup_{k\in\mathbb{N}}V_{\Gamma,k}\subseteq V_{\Gamma},\qquad\bigcup_{k\in\mathbb{N}}{\cal V}_{k}\subseteq{\cal V}$
are dense. To construct a sequence of approximate solutions, we now make the ansatz
[…]
for every $k\in\mathbb{N}$, where the coefficients $\mathbf{a}^{k}:=(a_{1}^{k},...,a_{k}^{k})^{\top}$, $\mathbf{b}^{k}:=(b_{1}^{k},...,b_{k}^{k})^{\top}$, $\mathbf{c}^{k}:=(c_{1}^{k},...,c_{k}^{k})^{\top}$, $\mathbf{d}^{k}:=(d_{1}^{k},...,d_{k}^{k})^{\top}$ are still to be determined.
Let now $k\in{\mathbb{N}}$ and $t\in[0,T]$ be arbitrary. We consider the bilinear form
$\mathcal{B}_{k,t}:{\bf V}_{\sigma,{\bf n}}\times{\bf V}_{\sigma,{\bf n}}\to{\mathbb{R}},$
$({\boldsymbol{v}},{\boldsymbol{w}})\mapsto 2\int_{\Omega}\nu\big{(}\varphi_{k}(t)\big{)}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda\big{(}\varphi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma\big{(}\psi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}},$
which is related to the weak formulation of the Brinkman equation with Navier-slip boundary condition. It is obvious that $\mathcal{B}_{k,t}$ is symmetric, and, in view of (A3), it is easy to see that $\mathcal{B}_{k,t}$ is continuous. We further recall that every ${\boldsymbol{v}}\in{\bf V}_{\sigma,{\bf n}}$ satisfies $\mathop{\rm div}\nolimits({\boldsymbol{v}})=0$ a.e. in $\Omega$, and ${\boldsymbol{v}}\cdot{\bf n}=0$ a.e. on $\Gamma$. By means of (A3), integration by parts and the Poincaré inequality presented in Lemma 2.1, we deduce
$\mathcal{B}_{k,t}({\boldsymbol{v}},{\boldsymbol{v}})$ $\geq 2\nu_{1}\int_{\Omega}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{v}}+\gamma_{1}\int_{\Gamma}|{\boldsymbol{v}}|^{2}$
$=\nu_{1}\int_{\Omega}|\nabla{\boldsymbol{v}}|^{2}+\gamma_{1}\int_{\Gamma}|{\boldsymbol{v}}|^{2}\geq C\|{\boldsymbol{v}}\|_{{\bf{V}}}^{2}$
for all ${\boldsymbol{v}}\in{\bf V}_{\sigma,{\bf n}}$. This means that the bilinear form $\mathcal{B}_{k,t}$ is coercive in ${\bf V}_{\sigma,{\bf n}}$. Hence, the Lax–Milgram lemma implies that there exists a unique function ${\boldsymbol{v}}_{k}(t)\in{\bf V}_{\sigma,{\bf n}}$ solving
$\mathcal{B}_{k,t}\big{(}{\boldsymbol{v}}_{k}(t),{\boldsymbol{w}}\big{)}=-\int_{\Omega}\varphi_{k}(t)\nabla\mu_{k}(t)\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi_{k}(t)\nabla_{\Gamma}\theta_{k}(t)\cdot{\boldsymbol{w}}$
for all ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$. As $t\in[0,T]$ was arbitrary, this defines a function ${\boldsymbol{v}}_{k}:[0,T]\to{\bf V}_{\sigma,{\bf n}}$. We point out that by construction, ${\boldsymbol{v}}_{k}$ depends continuously on the coefficients $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$.
We now want to adjust the coefficient vectors $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ such that the discretized weak formulation
$2\int_{\Omega}\nu(\varphi_{k})\mathbf{D}{\boldsymbol{v}}_{k}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda(\varphi_{k}){\boldsymbol{v}}_{k}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma(\psi_{k}){\boldsymbol{v}}_{k}\cdot{\boldsymbol{w}}$
$\quad=-\int_{\Omega}\varphi_{k}\nabla\mu_{k}\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi_{k}\nabla_{\Gamma}\theta_{k}\cdot{\boldsymbol{w}},$ (3.4a)
$\mathopen{\langle}\partial_{t}\varphi_{k},\zeta\mathclose{\rangle}_{V}-\int_{\Omega}\varphi_{k}{\boldsymbol{v}}_{k}\cdot\nabla\zeta+\int_{\Omega}M_{\Omega}(\varphi_{k})\nabla\mu_{k}\cdot\nabla\zeta=0,$ (3.4b)
$\mathopen{\langle}\partial_{t}\psi_{k},{\zeta_{\Gamma}}\mathclose{\rangle}_{{V}_{\Gamma}}-\int_{\Gamma}\psi_{k}{\boldsymbol{v}}_{k}\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}+\int_{\Gamma}M_{\Gamma}(\psi_{k})\nabla_{\Gamma}\theta_{k}\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}=0,$ (3.4c)
$\int_{\Omega}\mu_{k}\eta+\int_{\Gamma}\theta_{k}{\eta_{\Gamma}}=\int_{\Omega}\nabla\varphi_{k}\cdot\nabla\eta+\int_{\Omega}F^{\prime}(\varphi_{k})\eta+\int_{\Gamma}\nabla_{\Gamma}\psi_{k}\cdot\nabla_{\Gamma}{\eta_{\Gamma}}$
$\hskip 86.11084pt+\int_{\Gamma}G^{\prime}(\psi_{k}){\eta_{\Gamma}}+\frac{1}{K}\int_{\Gamma}(\psi_{k}-\varphi_{k})({\eta_{\Gamma}}-\eta)$ (3.4d)
for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V_{k}$, ${\zeta_{\Gamma}}\in V_{\Gamma,k}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}_{k}$, and the initial conditions
$\varphi_{k}(0)=\varphi_{0,k}:=\mathbb{P}_{V_{k}}(\varphi_{0})\quad\text{and}\quad\psi_{k}(0)=\psi_{0,k}:=\mathbb{P}_{V_{\Gamma,k}}(\psi_{0})$ (3.5)
are fulfilled. With the symbol $\mathbb{P}_{V_{k}}$ we denote the $H$-orthogonal projection of $V$ onto $V_{k}$ whereas $\mathbb{P}_{V_{\Gamma,k}}$ denotes the $H_{\Gamma}$-orthogonal projection of $V_{\Gamma}$ onto $V_{\Gamma,k}$.
Choosing $\zeta=u_{j}$ in (3.4b) and $\zeta_{\Gamma}=v_{j}$ in (3.4c) for $j=1,...,k$, we infer that $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is determined by a system of $2k$ nonlinear ordinary differential equations subject to the initial conditions
$[\mathbf{a}^{k}]_{i}(0)=a_{i}^{k}(0)=\left(\varphi_{0},u_{i}\right)_{H}\quad\text{and}\quad[\mathbf{b}^{k}]_{i}(0)=b_{i}^{k}(0)=\left(\psi_{0},v_{i}\right)_{H_{\Gamma}}$
for all $i\in\{1,...,k\}$. In particular, since the functions $M_{\Omega}$ and $M_{\Gamma}$ are continuous and ${\boldsymbol{v}}_{k}$ depends continuously on the coefficients $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$, the same holds for the right-hand side of this ODE system. Moreover, choosing $(\eta,\eta_{\Gamma})=(u_{j},0)$ and $(\eta,\eta_{\Gamma})=(0,v_{j})$ for $j=1,...,k$ in (3.1.1), respectively, we find that the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ are explicitly given by $2k$ algebraic equations whose right-hand side depends continuously on $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$. This allows us to replace $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ in the right-hand side of the aforementioned ODE system to obtain a closed ODE system for the vector-valued function $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ whose right-hand side depends continuously on $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$. Consequently, the Cauchy–Peano theorem implies the existence of at least one local-in-time solution
$(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}:[0,T_{k}^{*})\cap[0,T]\to{\mathbb{R}}^{2k}$
to the corresponding initial value problem. Here, we take $T_{k}^{*}>0$ as large as possible meaning that $[0,T_{k}^{*})\cap[0,T]$ is the right-maximal time interval of this solution. We can now reconstruct
$(\mathbf{c}^{k},\mathbf{d}^{k})^{\top}:[0,T_{k}^{*})\cap[0,T]\to{\mathbb{R}}^{2k}$
by the aforementioned system of $2k$ algebraic equations. Without loss of generality, we now assume $T_{k}^{*}\leq T$ to simplify the notation. Recalling the ansatz (3.3) as well as the construction of ${\boldsymbol{v}}_{k}$, we obtain an approximate solution $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ with
${\boldsymbol{v}}_{k}\in C^{1}([0,T_{k}^{*});{\bf V}_{\sigma,{\bf n}})\quad\text{and}\quad(\varphi_{k},\psi_{k}),(\mu_{k},\theta_{k})\in C^{1}([0,T_{k}^{*});{\cal V}),$
which fulfills the discretized weak formulation (3.1.1) on the time interval $[0,T_{k}^{*})$.
3.1.2   Uniform estimates
Let now $T_{k}\in(0,T_{k}^{*})$ be arbitrary. We derive suitable estimates for the approximate solutions $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ that are uniform in $k$ and $T_{k}$. These estimates will allow us to extend the approximate solutions onto the whole interval $[0,T]$ and to extract suitable convergent subsequences.
First estimate. We first test (3.4a) by ${\boldsymbol{v}}_{k}$, (3.4b) by $\mu_{k}$, (3.4c) by $\theta_{k}$, and (3.1.1) by $-(\partial_{t}\varphi_{k},\partial_{t}\psi_{k})$. We then add these equations and integrate the resulting equation with respect to time from 00 to an arbitrary $t\in[0,T_{k}]$. We obtain
$E_{K}\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}+2\int_{0}^{t}\int_{\Omega}\nu(\varphi_{k})\left|\mathbf{D}{\boldsymbol{v}}_{k}\right|^{2}+\int_{0}^{t}\int_{\Omega}\lambda(\varphi_{k})\left|{\boldsymbol{v}}_{k}\right|^{2}+\int_{0}^{t}\int_{\Gamma}\gamma(\psi_{k})\left|{\boldsymbol{v}}_{k}\right|^{2}$
$\quad+\int_{0}^{t}\int_{\Omega}M_{\Omega}(\varphi_{k})\left|\nabla\mu_{k}\right|^{2}+\int_{0}^{t}\int_{\Gamma}M_{\Gamma}(\psi_{k})\left|\nabla_{\Gamma}\theta_{k}\right|^{2}\leq E_{K}(\varphi_{0,k},\psi_{0,k})$ (3.6)
for every $t\in[0,T_{k}]$. Due to (3.5) and the assumptions on the initial data, we have
$\left\|\varphi_{0,k}\right\|_{V}\leq C\left\|\varphi_{0}\right\|_{V}\leq C\quad\text{and}\quad\left\|\psi_{0,k}\right\|_{{V}_{\Gamma}}\leq C\left\|\psi_{0}\right\|_{{V}_{\Gamma}}\leq C.$ (3.7)
In view of the growth conditions from (R1), this directly implies
$\left\|F\big{(}\varphi_{0,k}\big{)}\right\|_{L^{1}(\Omega)}\leq C,\quad\left\|G\big{(}\psi_{0,k}\big{)}\right\|_{L^{1}(\Gamma)}\leq C\quad\text{and thus},\quad E_{K}(\varphi_{0,k},\psi_{0,k})\leq C.$ (3.8)
Hence, using the conditions in (A2) and (A3), a straightforward computation yields
$\left\|\mathbf{D}{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\mathbb{H}})}+\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};\mathbf{H}_{\Gamma})}+\left\|\nabla\varphi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}})}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}}_{\Gamma})}$
$\quad+\left\|\nabla\mu_{k}\right\|_{L^{2}(0,T_{k};{\bf{H}})}+\left\|\nabla_{\Gamma}\theta_{k}\right\|_{L^{2}(0,T_{k};{\bf{H}}_{\Gamma})}\leq C.$ (3.9)
Invoking the Poincaré inequality presented in Lemma 2.1, we directly infer
$\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\bf V}_{\sigma,{\bf n}})}\leq C.$ (3.10)
Next, taking $\zeta=\frac{1}{|\Omega|}$ in (3.4b), and $\zeta_{\Gamma}=\frac{1}{|\Gamma|}$ in (3.4c), we infer
$\mathopen{\langle}\varphi_{k}(t)\mathclose{\rangle}_{\Omega}=\mathopen{\langle}\varphi_{k,0}\mathclose{\rangle}_{\Omega},\quad\mathopen{\langle}\psi_{k}(t)\mathclose{\rangle}_{\Gamma}=\mathopen{\langle}\psi_{{k,0}}\mathclose{\rangle}_{\Gamma}\quad\text{for all $t\in[0,T_{k}]$}.$
Hence, in view of (3.7) and (3.1.2), we use the Poincaré–Wirtinger inequality to conclude
$\left\|\varphi_{k}\right\|_{L^{\infty}(0,T_{k};V)}+\left\|\psi_{k}\right\|_{L^{\infty}(0,T_{k};{V}_{\Gamma})}\leq C.$ (3.11)
Second estimate. Let now $\zeta\in L^{2}(0,T_{k};V)$ and ${\zeta_{\Gamma}}\in L^{2}(0,T_{k};{V}_{\Gamma})$ be arbitrary test functions. Testing (3.4b) with ${\overline{\zeta}}:=\mathbb{P}_{V_{k}}(\zeta)$ and exploiting (3.1.2)–(3.11) along with Sobolev’s embeddings, we obtain
$\left|\int_{0}^{T_{k}}\mathopen{\langle}\partial_{t}\varphi_{k},\zeta\mathclose{\rangle}_{V}\;\right|=\left|\int_{0}^{T_{k}}\mathopen{\langle}\partial_{t}\varphi_{k},{\overline{\zeta}}\mathclose{\rangle}_{V}\;\right|$
$\quad=\left|\int_{0}^{T_{k}}\!\!\int_{\Omega}\varphi_{k}{\boldsymbol{v}}_{k}\cdot\nabla{\overline{\zeta}}-\int_{0}^{T_{k}}\!\!\int_{\Omega}M_{\Omega}(\varphi_{k})\nabla\mu_{k}\cdot\nabla{\overline{\zeta}}\;\right|$
$\quad\leq\left(\left\|\varphi_{k}\right\|_{L^{\infty}(0,T_{k};L^{4}(\Omega))}\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};(L^{4}(\Omega))^{d})}+M_{2}\left\|\nabla\mu_{k}\right\|_{L^{2}(0,T_{k};{\bf H})}\right)\left\|{\overline{\zeta}}\right\|_{L^{2}(0,T_{k};V)}$
$\quad\leq C\left\|\zeta\right\|_{L^{2}(0,T_{k};V)}.$ (3.12)
Hence, taking the supremum over all $\zeta\in L^{2}(0,T_{k};V)$ with $\left\|\zeta\right\|_{L^{2}(0,T_{k};V)}\leq 1$, we deduce
$\left\|\partial_{t}\varphi_{k}\right\|_{L^{2}(0,T_{k};V^{*})}\leq C.$ (3.13)
Proceeding similarly and testing (3.4c) with ${{\overline{\zeta}}_{\Gamma}}:=\mathbb{P}_{V_{\Gamma,k}}({\zeta_{\Gamma}})$, we obtain the estimate
$\left|\int_{0}^{T_{k}}\mathopen{\langle}\partial_{t}\psi_{k},{\zeta_{\Gamma}}\mathclose{\rangle}_{{V}_{\Gamma}}\;\right|$
$\leq\Bigl{(}C\left\|\psi_{k}\right\|_{L^{\infty}(0,T_{k};{V}_{\Gamma})}\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\bf V}_{\sigma,{\bf n}})}+M_{\Gamma,2}\left\|\nabla_{\Gamma}\theta_{k}\right\|_{L^{2}(0,T_{k};{\bf H}_{\Gamma})}\Bigr{)}\left\|{{\overline{\zeta}}}_{\Gamma}\right\|_{L^{2}(0,T_{k};{V}_{\Gamma})}$
$\leq C\left\|{\zeta_{\Gamma}}\right\|_{L^{2}(0,T_{k};{V}_{\Gamma})}.$ (3.14)
Taking the supremum over all ${\zeta_{\Gamma}}\in L^{2}(0,T_{k};{V}_{\Gamma})$ with $\left\|{\zeta_{\Gamma}}\right\|_{L^{2}(0,T_{k};{V}_{\Gamma})}\leq 1$, we conclude
$\left\|\partial_{t}\psi_{k}\right\|_{L^{2}(0,T_{k};V_{\Gamma}^{*})}\leq C.$ (3.15)
Third estimate. Next, we want to derive uniform bounds on $\mu_{k}$ in $L^{4}(0,T_{k};H)\cap L^{2}(0,T_{k};V)$ and on $\theta_{k}$ in $L^{4}(0,T_{k};{H}_{\Gamma})\cap L^{2}(0,T_{k};{V}_{\Gamma})$. Therefore, we choose arbitrary functions $\eta\in L^{1}(0,T_{k};V)$ and $\eta_{\Gamma}\in L^{1}(0,T_{k};{V}_{\Gamma})$ and we set ${\overline{\eta}}:=\mathbb{P}_{V_{k}}(\eta)$ and ${\overline{\eta_{\Gamma}}}:=\mathbb{P}_{V_{\Gamma,k}}(\eta_{\Gamma})$. Testing (3.1.1) by $({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})$, recalling the growth conditions from (R1) as well as the uniform bounds (3.1.2) and (3.10), we use Hölder’s inequality and Sobolev’s embedding theorem to derive the estimate
$\int_{0}^{T_{k}}\big{|}\mathopen{\langle}(\mu_{k},\theta_{k}),(\eta,{\eta_{\Gamma}})\mathclose{\rangle}_{{\cal V}}\big{|}=\int_{0}^{T_{k}}\big{|}\mathopen{\langle}(\mu_{k},\theta_{k}),({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})\mathclose{\rangle}_{{\cal V}}\big{|}$
$\quad\leq\int_{0}^{T_{k}}\Big{[}\left\|\nabla\varphi_{k}\right\|_{{\bf H}}\left\|\nabla{\overline{\eta}}\right\|_{{\bf H}}+\left\|F^{\prime}(\varphi_{k})\right\|_{L^{\frac{6}{5}}}(\Omega)\left\|{\overline{\eta}}\right\|_{L^{6}(\Omega)}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{{\bf H}_{\Gamma}}\left\|\nabla_{\Gamma}{\overline{{\eta_{\Gamma}}}}\right\|_{{\bf H}_{\Gamma}}$
$\qquad\qquad\quad+\left\|G^{\prime}(\psi_{k})\right\|_{{H}_{\Gamma}}\left\|{\overline{{\eta_{\Gamma}}}}\right\|_{{H}_{\Gamma}}+\tfrac{1}{K}\left\|\psi_{k}-\varphi_{k}\right\|_{{H}_{\Gamma}}\left\|{\overline{{\eta_{\Gamma}}}}-{\overline{\eta}}\right\|_{{H}_{\Gamma}}\Big{]}$
$\quad\leq C(1+\left\|\varphi_{k}\right\|_{L^{\infty}(0,T_{k};V)}^{p-1}+\left\|\psi_{k}\right\|_{L^{\infty}(0,T_{k};{V}_{\Gamma})}^{q-1})\left\|(\eta,{\eta_{\Gamma}})\right\|_{L^{1}(0,T_{k};{\cal V})}$
in $[0,T_{k}]$. Taking the supremum over all $(\eta,{\eta_{\Gamma}})\in L^{1}(0,T_{k};{\cal V})$ with $\left\|(\eta,{\eta_{\Gamma}})\right\|_{L^{1}(0,T_{k};{\cal V})}\leq 1$, and using (3.11), we infer
$\left\|(\mu_{k},\theta_{k})\right\|_{L^{\infty}(0,T_{k};{\cal V}^{*})}\leq C.$ (3.16)
We further have
$\left\|(\mu_{k},\theta_{k})\right\|^{2}_{\cal H}$ $=\mathopen{\langle}(\mu_{k},\theta_{k}),(\mu_{k},\theta_{k})\mathclose{\rangle}_{{\cal V}}$
$\leq C\left\|(\mu_{k},\theta_{k})\right\|_{{{\cal V}^{*}}}\big{(}\left\|(\mu_{k},\theta_{k})\right\|_{\cal H}+\left\|(\nabla\mu_{k},\nabla_{\Gamma}\theta_{k})\right\|_{{\bf H}\times{\bf H}_{\Gamma}}\big{)}$
$\leq\frac{1}{2}\left\|(\mu_{k},\theta_{k})\right\|^{2}_{\cal H}+C\left\|(\mu_{k},\theta_{k})\right\|_{{{\cal V}^{*}}}\left\|(\nabla\mu_{k},\nabla_{\Gamma}\theta_{k})\right\|_{{\bf H}\times{\bf H}_{\Gamma}}+C\left\|(\mu_{k},\theta_{k})\right\|_{{{\cal V}^{*}}}^{2}.$
Hence, squaring and integrating this estimate with respect to time, we use (3.1.2) and (3.16) to conclude
$\left\|(\mu_{k},\theta_{k})\right\|_{L^{4}(0,T_{k};{\cal H})}\leq C.$ (3.17)
In particular, we thus have
$\left\|\mu_{k}\right\|_{L^{2}(0,T_{k};V)}+\left\|\theta_{k}\right\|_{L^{2}(0,T_{k};{V}_{\Gamma})}\leq C.$ (3.18)
Overall estimate. Combining (3.1.2)–(3.11), (3.13), (3.15), (3.17) and (3.18), we obtain the overall uniform estimate
$\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\bf V}_{\sigma,{\bf n}})\cap L^{2}(0,T_{k};\mathbf{H}_{\Gamma})}+\left\|\varphi_{k}\right\|_{H^{1}(0,T_{k};{V^{*}})\cap L^{\infty}(0,T_{k};V)}+\left\|\psi_{k}\right\|_{H^{1}(0,T_{k};{{V}_{\Gamma}^{*}})\cap L^{\infty}(0,T_{k};{V}_{\Gamma})}$
$\quad+\left\|\mu_{k}\right\|_{L^{4}(0,T_{k};H)\cap L^{2}(0,T_{k};V)}+\left\|\theta_{k}\right\|_{L^{4}(0,T_{k};{H}_{\Gamma})\cap L^{2}(0,T_{k};{V}_{\Gamma})}\leq C.$ (3.19)
3.1.3   Extension of the approximate solution onto the whole time interval $[0,T]$
In Step 1, we constructed the coefficients $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ as a solution of a nonlinear system of ODEs existing on its right-maximal time interval $[0,T_{k}^{*})\cap[0,T]$. We now assume that $T_{k}^{*}\leq T$. By the definition of the approximate solutions given in (3.3) and the uniform bound (3.1.2), we infer that for any $T_{k}\in[0,T_{k}^{*})$, all $t\in[0,T_{k}]$, and all $i\in\{1,...,k\}$,
$|a_{i}^{k}(t)|+|b_{i}^{k}(t)|$ $=\big{|}\left(\varphi_{k}(t),u_{i}\right)_{H}\big{|}+\big{|}\left(\psi_{k}(t),v_{i}\right)_{{H}_{\Gamma}}\big{|}$
$\leq\|\varphi_{k}\|_{L^{\infty}(0,T_{k};H)}+\|\psi_{k}\|_{L^{\infty}(0,T_{k};{H}_{\Gamma})}\leq C.$
This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is bounded on the time interval $[0,T_{k}^{*})$ by a constant that is independent of $T_{k}$ and $k$. Hence, according to classical ODE theory, the solution can thus be extended beyond the time $T_{k}^{*}$. However, as the solution was assumed to be right-maximal, this is a contradiction. We thus have $T_{k}^{*}>T$, which directly implies $[0,T_{k}^{*})\cap[0,T]=[0,T]$. This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ of the ODE system actually exists on the whole time interval $[0,T]$. As the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ can be reconstructed from $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$ by the corresponding system of algebraic equations, they also exist on the whole time interval $[0,T]$. Recalling (3.3) and the construction of ${\boldsymbol{v}}_{k}$ this directly entails that the approximate solution $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ actually exists in $[0,T]$. Hence, choosing $T_{k}=T$ in (3.1.2), we eventually conclude
$\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})\cap L^{2}(0,T;\mathbf{H}_{\Gamma})}+\left\|\varphi_{k}\right\|_{H^{1}(0,T;V^{*})\cap L^{\infty}(0,T;V)}+\left\|\psi_{k}\right\|_{H^{1}(0,T;{V}_{\Gamma}^{*})\cap L^{\infty}(0,T;{V}_{\Gamma})}$
$\quad+\left\|\mu_{k}\right\|_{L^{4}(0,T;H)\cap L^{2}(0,T;V)}+\left\|\theta_{k}\right\|_{L^{4}(0,T;{H}_{\Gamma})\cap L^{2}(0,T;{V}_{\Gamma})}\leq C.$ (3.20)
3.1.4   Convergence to a weak solution as $k\to\infty$
Considering the uniform estimate (3.1.3), we use the Banach–Alaoglu theorem and the Aubin–Lions–Simon lemma to infer that there exist functions ${\boldsymbol{v}}$, $\varphi$, $\mu$, $\psi$ and $\theta$ such that for any $s\in[0,1)$,
${\boldsymbol{v}}_{k}$ $\to{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (3.21a)
${\boldsymbol{v}}_{k}|_{\Gamma}$ $\to{\boldsymbol{v}}|_{\Gamma}$ weakly in $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (3.21b)
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$ (3.21c)
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$ (3.21d)
$\mu_{k}$ $\to\mu$ weakly in $\quad\text{weakly in $L^{2}(0,T;V)\cap L^{4}(0,T;H)$},$ (3.21e)
$\theta_{k}$ $\to\theta$ weakly in $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})\cap L^{4}(0,T;{H}_{\Gamma})$},$ (3.21f)
as $k\to\infty$ along a non-relabeled subsequence. In particular, this shows that the functions ${\boldsymbol{v}}$, $\varphi$, $\psi$, $\mu$ and $\theta$ have the regularity demanded in Definition 2.3(i).
Due to the trace theorem, the strong convergence from (3.1.4) (with $s>\frac{1}{2}$) directly yields
$\varphi_{k}|_{\Gamma}\to\varphi|_{\Gamma}\quad\text{strongly in $C^{0}([0,T];{H}_{\Gamma})$}.$ (3.22)
Recalling the growth conditions from (R1), we further deduce from the uniform bound (3.1.3) that
$\left\|F^{\prime}(\varphi_{k})\right\|_{L^{\frac{6}{5}}(Q)}\leq C\quad\text{and}\quad\left\|G^{\prime}(\varphi_{k})\right\|_{L^{2}(\Sigma)}\leq C.$
Hence, there exist weakly convergent subsequences in the respective spaces. As $F^{\prime}$ and $G^{\prime}$ are continuous, we use the pointwise convergences from (3.1.4) and (3.1.4) to conclude
$F^{\prime}(\varphi_{k})$ $\to F^{\prime}(\varphi)$ weakly in L65(Q) and a.e. in $\quad\text{weakly in $L^{\frac{6}{5}}(Q)$ and {a.e.\ in~{}$Q$}},$ (3.23)
$G^{\prime}(\psi_{k})$ $\to G^{\prime}(\psi)$ weakly in $L^{2}(\Sigma)$ and a.e. on $\Sigma$ (3.24)
since the weak limit and the pointwise limit must coincide (see, e.g., [26, Proposition 9.2c]). Furthermore, it follows from the pointwise convergences in (3.1.4) and (3.1.4) that, as $k\to\infty$,
[…]
$M_{\Gamma}(\psi_{k})$ $\to M_{\Gamma}(\psi),$ $\quad\gamma(\psi_{k})$ $\to\gamma(\psi)$ a.e. on $\Sigma$ (3.26)
as the functions $M_{\Omega}$, $M_{\Gamma}$, $\nu$, $\lambda$ and $\gamma$ are continuous. Since, due to (A2) and (A3), these functions are also bounded, we use Lebesgue’s dominated convergence theorem along with the weak convergences in (3.1.4) to infer that
$\nu(\varphi_{k})\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\nu(\varphi)\mathbf{D}{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.27)
$\lambda(\varphi_{k}){\boldsymbol{v}}_{k}$ $\to\lambda(\varphi){\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.28)
$M_{\Omega}(\varphi_{k})\nabla\mu_{k}$ $\to M_{\Omega}(\varphi)\nabla\mu$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.29)
$\gamma(\psi_{k}){\boldsymbol{v}}_{k}$ $\to\gamma(\psi){\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.30)
$M_{\Gamma}(\psi_{k})\nabla_{\Gamma}\theta_{k}$ $\to M_{\Gamma}(\psi)\nabla_{\Gamma}\theta$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.31)
Combining the convergences (3.21a)–(3.21f), (3.22)–(3.24) and (3.27)–(3.31), it is straightforward to pass to the limit as $k\to\infty$ in the discretized weak formulation (3.1.1) to conclude that the quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ fulfills the variational formulation ((ii)) for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$,
$\zeta\in\bigcup_{k\in\mathbb{N}}V_{k}\subseteq V,\qquad{\zeta_{\Gamma}}\in\bigcup_{k\in\mathbb{N}}V_{\Gamma,k}\subseteq V_{\Gamma},\qquad(\eta,{\eta_{\Gamma}})\in\bigcup_{k\in\mathbb{N}}{\cal V}_{k}\subseteq{\cal V}.$
Hence, because of density, ((ii)) holds true for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, ${\zeta_{\Gamma}}\in V_{\Gamma}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}={\cal V}_{K}$. This verifies Definition 2.3(ii).
Moreover, we deduce from (3.5) that
$\big{(}\varphi_{k}(0),\psi_{k}(0)\big{)}\to\big{(}\varphi_{0},\psi_{0}\big{)}\quad\text{strongly in ${\cal H}$}$
as the orthogonal projections converge strongly in $H$ and in ${H}_{\Gamma}$, respectively. On the other hand, it follows from the strong convergences in (3.1.4) and (3.1.4) that
$\big{(}\varphi_{k}(0),\psi_{k}(0)\big{)}\to\big{(}\varphi(0),\psi(0)\big{)}\quad\text{strongly in ${\cal H}$.}$
Hence, due to the uniqueness of the limit, this verifies Definition 2.3(iii).
We still need to establish the weak energy dissipation law. Therefore, let $\rho\in C^{\infty}([0,T])$ be an arbitrary nonnegative test function. Employing the convergences (3.1.4) and (3.1.4), the weak lower semicontinuity of the mappings
$L^{2}(0,T;V)$ $\ni\zeta\mapsto\int_{0}^{T}\left\|\nabla\zeta(t)\right\|_{{\bf H}}^{2}\rho(t),$
$L^{2}(0,T;{V}_{\Gamma})$ $\ni\xi\mapsto\int_{0}^{T}\left\|\nabla_{\Gamma}\xi(t)\right\|_{{\bf H}_{\Gamma}}^{2}\rho(t),$
as well as Fatou’s lemma, we deduce
$\int_{0}^{T}E_{K}\big{(}\varphi(t),\psi(t)\big{)}\,\rho(t)\leq\underset{k\to\infty}{\lim\inf}\;\int_{0}^{T}E_{K}\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\,\rho(t).$ (3.32)
Proceeding similarly as above, we derive the convergences
$\sqrt{\nu(\varphi_{k})}\,\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\sqrt{\nu(\varphi)}\,\mathbf{D}{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.33)
$\sqrt{\lambda(\varphi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\lambda(\varphi)}\,{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.34)
$\sqrt{M_{\Omega}(\varphi_{k})}\,\nabla\mu_{k}$ $\to\sqrt{M_{\Omega}(\varphi)}\,\nabla\mu$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.35)
$\sqrt{\gamma(\psi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\gamma(\psi)}\,{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.36)
$\sqrt{M_{\Gamma}(\psi_{k})}\,\nabla_{\Gamma}\theta_{k}$ $\to\sqrt{M_{\Gamma}(\psi)}\,\nabla_{\Gamma}\theta$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.37)
Hence, employing (3.1.2), (3.32) and weak lower semicontinuity, we eventually obtain
$\int_{0}^{T}E_{K}\big{(}\varphi(t),\psi(t)\big{)}\rho(t)$
$\qquad+\int_{0}^{T}\int_{\Omega}\Big{[}2\nu(\varphi)\left|\mathbf{D}{\boldsymbol{v}}\right|^{2}\rho(t)+\lambda(\varphi)\left|{\boldsymbol{v}}\right|^{2}\rho(t)+M_{\Omega}(\varphi)\left|\nabla\mu\right|^{2}\rho(t)\Big{]}$
$\qquad+\int_{0}^{T}\int_{\Gamma}\Big{[}\gamma(\psi)\left|{\boldsymbol{v}}\right|^{2}\rho(t)+M_{\Gamma}(\psi)\left|\nabla_{\Gamma}\theta\right|^{2}\rho(t)\Big{]}$
$\quad\leq\underset{k\to\infty}{\liminf}\Bigg{\{}\int_{0}^{T}E_{K}\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\rho(t)$
$\qquad+\int_{0}^{T}\int_{\Omega}\Big{[}2\nu(\varphi_{k})\left|\mathbf{D}{\boldsymbol{v}}_{k}\right|^{2}\rho(t)+\lambda(\varphi_{k})\left|{\boldsymbol{v}}_{k}\right|^{2}\rho(t)+M_{\Omega}(\varphi_{k})\left|\nabla\mu_{k}\right|^{2}\rho(t)\Big{]}$
$\qquad+\int_{0}^{T}\int_{\Gamma}\Big{[}\gamma(\psi_{k})\left|{\boldsymbol{v}}_{k}\right|^{2}\rho(t)+M_{\Gamma}(\psi_{k})\left|\nabla_{\Gamma}\theta_{k}\right|^{2}\rho(t)\Big{]}\Bigg{\}}$
$\quad\leq\underset{k\to\infty}{\lim}\int_{0}^{T}E_{K}\big{(}\varphi_{0,k},\psi_{0,k}\big{)}\rho(t)=\int_{0}^{T}E_{K}\big{(}\varphi_{0},\psi_{0}\big{)}\rho(t).$
Here, invoking the growth conditions from (R1), the equality in the last line follows by means of Lebesgue’s general convergence theorem (see [8, Section 3.25]) since the orthogonal projections in (3.5) converge strongly in $V$ and in ${V}_{\Gamma}$, respectively. As the nonnegative test function $\rho$ was arbitrary, this proves that the weak energy dissipation law stated in ((iv)) holds for almost all $t\in[0,T]$. As the time integral in this inequality is continuous with respect to $t$ and since the mapping $t\mapsto E_{K}\big{(}\varphi(t),\psi(t)\big{)}$ is lower semicontinuous, we conclude that ((iv)) actually holds true for all $t\in[0,T]$. This means that Definition 2.3(iv) is verified.
We have thus shown that the quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ is a weak solution in the sense of Definition 2.3.
3.1.5   Additional regularity for the phase-fields
To prove additional regularity for the phase-fields, we assume $\Omega$ to be of class $C^{\ell}$ with $\ell\in\{2,3\}$, and we return to the Faedo–Galerkin scheme. First of all, applying regularity theory for Poisson’s equation with an inhomogeneous Neumann boundary condition (see, e.g., [52, Theorem 4.18] or [59, s. 5, Proposition 7.7]), we infer that the eigenfunctions $u_{i}$ exhibit the regularity $u_{i}\in H^{\ell}(\Omega)$ for all $i\in{\mathbb{N}}$. Moreover, applying regularity theory for elliptic equations on submanifolds (see, e.g., [59, s. 5, Theorem 1.3]), we conclude that the eigenfunctions $v_{i}$ have the regularity $v_{i}\in H^{\ell}(\Gamma)$ for all $i\in{\mathbb{N}}$. This directly entails
$\varphi_{k}(t)\in V_{k}\cap H^{\ell}(\Omega)\quad\text{and}\quad\psi_{k}(t)\in V_{\Gamma,k}\cap H^{\ell}(\Gamma)$
for all $t\in[0,T]$.
Let now $t\in[0,T]$ be arbitrary. We infer from (3.1.1) that there exists a null set $\mathcal{N}\subset[0,T]$ such that
$\int_{\Omega}\nabla\varphi_{k}(t)\cdot\nabla\eta+\int_{\Gamma}\nabla_{\Gamma}\psi_{k}(t)\cdot\nabla_{\Gamma}{\eta_{\Gamma}}+\frac{1}{K}\int_{\Gamma}\big{(}\psi_{k}(t)-\varphi_{k}(t)\big{)}({\eta_{\Gamma}}-\eta)$
$\quad=\int_{\Omega}\Big{(}\mu_{k}(t)-\mathbb{P}_{V_{k}}\big{[}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{]}\Big{)}\,\eta+\int_{\Gamma}\Big{(}\theta_{k}(t)-\mathbb{P}_{V_{\Gamma,k}}\big{[}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{]}\Big{)}\,{\eta_{\Gamma}}$ (3.38)
for all test functions $(\eta,{\eta_{\Gamma}})\in{\cal V}$ and all $t\in[0,T]\setminus\mathcal{N}$.
Let now $t\in[0,T]\setminus\mathcal{N}$ be arbitrary. Then, (3.1.5) entails that the pair $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}$ is a weak solution of the bulk-surface elliptic problem
$-\Delta\varphi_{k}(t)$ $=f_{k}(t)$ in $\quad\text{in $\Omega$},$ (3.39a)
$-\Delta_{\Gamma}\psi_{k}(t)+\partial_{\bf n}\varphi_{k}(t)$ $=g_{k}(t)$ on $\quad\text{on $\Gamma$},$ (3.39b)
$K\partial_{\bf n}\varphi_{k}(t)$ $=\psi_{k}(t)-\varphi_{k}(t)$ on $\quad\text{on $\Gamma$},$ (3.39c)
where
$f_{k}(t)$ $:=\mu_{k}(t)-\mathbb{P}_{V_{k}}\big{[}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{]}$ $\in V_{k}\subset V,$
$g_{k}(t)$ $:=\theta_{k}(t)-\mathbb{P}_{V_{\Gamma,k}}\big{[}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{]}$ $\in V_{\Gamma,k}\subset{V}_{\Gamma}.$
In the following, without loss of generality, we assume that the growth conditions from (R1) hold with $p\in[5,6)$. Most part of the computations to follow resemble those performed in [40]. Therefore, we just highlight the key steps for the reader’s convenience. Using the uniform estimate (3.1.3), the Gagliardo–Nirenberg inequality and Young’s inequality, we derive the estimate
$\big{\|}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{\|}_{H}$ $\leq C+C\|\varphi_{k}(t)\|_{L^{2(p-1)}(\Omega)}^{p-1}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{6}(\Omega)}^{\frac{p+2}{2}}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-4}{2}}$
$\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{2}(\Omega)},$ (3.40)
for any $\varepsilon\in(0,1)$. Moreover, due to (R1) and (3.1.3), we have
$\big{\|}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{\|}_{{H}_{\Gamma}}\leq C+C\|\psi_{k}(t)\|_{L^{2(q-1)}(\Gamma)}^{q-1}\leq C.$ (3.41)
We first consider the case $\ell=2$. Applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we deduce $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
$\left\|\varphi_{k}(t)\right\|_{H^{2}(\Omega)}^{2}+\left\|\psi_{k}(t)\right\|_{H^{2}(\Gamma)}^{2}\leq C\left\|f_{k}(t)\right\|_{H}^{2}+C\left\|g_{k}(t)\right\|_{{H}_{\Gamma}}^{2}$
$\quad\leq C\left\|\mu_{k}(t)\right\|_{H}^{2}+C\left\|\theta_{k}(t)\right\|_{{H}_{\Gamma}}^{2}+C(1+\varepsilon^{-2})+C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{2}.$
After choosing $\varepsilon\in(0,1)$ sufficiently small, we absorb the term $C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{2}$ by the left-hand side. Squaring and integrating the resulting inequality with respect to $t$ from 00 to $T$, we use (3.1.3) to arrive at the uniform estimate
$\left\|\varphi_{k}\right\|_{L^{4}(0,T;H^{2}(\Omega))}+\left\|\psi_{k}\right\|_{L^{4}(0,T;H^{2}(\Gamma))}\leq C.$ (3.42)
Consequently, we have, as $k\to\infty$,
$\varphi_{k}$ $\to\varphi\quad\text{weakly in $L^{4}(0,T;H^{2}(\Omega))$},$
$\psi_{k}$ $\to\psi\quad\text{weakly in $L^{4}(0,T;H^{2}(\Gamma))$},$
after another subsequence extraction. This proves the regularity assertion (2.15a). As a direct consequence, it follows via integration by parts and the fundamental lemma of calculus that the equations (1.1d), (1.1f) and (1.1g) are now even satisfied in the strong sense (i.e., a.e. in $Q$ and on $\Sigma$, respectively).
Let us now consider the case $\ell=3$. Employing the uniform estimate (3.1.3), the Gagliardo–Nirenberg inequality, Agmon’s inequality and Young’s inequality, we derive the estimate
$\big{\|}F^{\prime\prime}\big{(}\varphi_{k}(t)\big{)}\nabla\varphi_{k}(t)\big{\|}_{{\bf H}}$ $\leq C\|\nabla\varphi_{k}(t)\|_{{\bf H}}+C\||\varphi_{k}(t)|^{p-2}\nabla\varphi_{k}(t)\|_{{\bf H}}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{2(p-2)}(\Omega)}^{p-2}\|\nabla\varphi_{k}(t)\|_{{\bf{L}}^{\infty}(\Omega)}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{6}(\Omega)}^{\frac{p+1}{2}}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-5}{2}+\frac{1}{2}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-4}{2}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{1}(\Omega)}^{\frac{p-4}{4}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{p-4}{4}+\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{p-2}{4}}$
$\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{3}(\Omega)}$ (3.43)
for any $\varepsilon\in(0,1)$. In combination with (3.1.5), this proves that
$\big{\|}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{\|}_{V}\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{3}(\Omega)}.$ (3.44)
In a similar fashion, we derive the estimate
$\big{\|}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{\|}_{{V}_{\Gamma}}\leq C\varepsilon^{-1}+\varepsilon\|\psi_{k}(t)\|_{H^{3}(\Gamma)}$ (3.45)
for any $\varepsilon\in(0,1)$. Applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we infer $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\in H^{3}(\Omega)\times H^{3}(\Gamma)$ with
$\left\|\varphi_{k}(t)\right\|_{H^{3}(\Omega)}^{2}+\left\|\psi_{k}(t)\right\|_{H^{3}(\Gamma)}^{2}\leq C\left\|f_{k}(t)\right\|_{V}^{2}+C\left\|g_{k}(t)\right\|_{{V}_{\Gamma}}^{2}$
$\quad\leq C\left\|\mu_{k}(t)\right\|_{V}^{2}+C\left\|\theta_{k}(t)\right\|_{{V}_{\Gamma}}^{2}+C\varepsilon^{-2}+C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{2}+C\varepsilon^{2}\|\psi_{k}(t)\|_{H^{3}(\Gamma)}^{2}.$
After choosing $\varepsilon\in(0,1)$ sufficiently small, we absorb the terms $C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{2}$ and $C\varepsilon^{2}\|\psi_{k}(t)\|_{H^{3}(\Gamma)}^{2}$ by the left-hand side. Integrating the resulting estimate with respect to $t$ from 00 to $T$, we use (3.1.3) to eventually obtain the uniform estimate
$\left\|\varphi_{k}\right\|_{L^{2}(0,T;H^{3}(\Omega))}+\left\|\psi_{k}\right\|_{L^{2}(0,T;H^{3}(\Gamma))}\leq C.$ (3.46)
Therefore, after another subsequence extraction, we have, as $k\to\infty$,
$\varphi_{k}$ $\to\varphi\quad\text{weakly in $L^{2}(0,T;H^{3}(\Omega))$},$
$\psi_{k}$ $\to\psi\quad\text{weakly in $L^{2}(0,T;H^{3}(\Gamma))$}.$
Hence, the regularity assertion (2.15b) is verified, and the regularity (2.16) follows directly from Proposition A.1.
This means that all claims are established and thus, the proof is complete. ∎
3.2   The limit $K\to 0$ and existence of a weak solution in the case $K=0$
Proof of Theorem 2.5.
In this proof, the letter $C$ will denote generic positive constants that may depend on $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3), but not on ${K_{n}}$ or $n$. Such constants may also change their value from line to line.
First of all, as the initial data were prescribed as $(\varphi_{0},\psi_{0})\in{\cal V}_{0}$, they satisfy the Dirichlet type coupling condition $\varphi_{0}|_{\Gamma}=\psi_{0}$ a.e. on $\Gamma$. In view of the definition of the energy functional in (1), this means that the ${K_{n}}$-depending term in the energy $E_{K_{n}}(\varphi_{0},\psi_{0})$ vanishes. It thus holds
$E_{K_{n}}(\varphi_{0},\psi_{0})=E_{0}(\varphi_{0},\psi_{0})\leq C\quad\text{for all $n\in{\mathbb{N}}$}.$ (3.47)
According to Definition 2.3(iv), the solutions $({\boldsymbol{v}}^{K_{n}},\varphi^{K_{n}},\mu^{K_{n}},\psi^{K_{n}},\theta^{K_{n}})$ satisfy the weak energy dissipation law. By the definition of $E_{K_{n}}$, we have
$E_{K_{n}}\big{(}\varphi^{K_{n}}(t),\psi^{K_{n}}(t)\big{)}+2\int_{0}^{t}\int_{\Omega}\nu(\varphi^{K_{n}})|\mathbf{D}{\boldsymbol{v}}^{K_{n}}|^{2}+\int_{0}^{t}\int_{\Omega}\lambda(\varphi^{K_{n}})|{\boldsymbol{v}}^{K_{n}}|^{2}$
$\qquad+\int_{0}^{t}\int_{\Gamma}\gamma(\psi^{K_{n}})|{\boldsymbol{v}}^{K_{n}}|^{2}+\int_{0}^{t}\int_{\Omega}M_{\Omega}(\varphi^{K_{n}})|\nabla\mu^{K_{n}}|^{2}+\int_{0}^{t}\int_{\Gamma}M_{\Gamma}(\psi^{K_{n}})|\nabla_{\Gamma}\theta^{K_{n}}|^{2}$
$\quad\leq E_{K_{n}}(\varphi_{0},\psi_{0})\leq C$
for all $t\in[0,T]$ and all $n\in{{\mathbb{N}}}$. In particular, recalling that the potentials $F$ and $G$ are nonnegative, this directly yields
$\left\|\varphi^{K_{n}}-\psi^{K_{n}}\right\|_{{H}_{\Gamma}}^{2}\leq C{K_{n}}\quad\text{for all $t\in[0,T]$ and all $n\in{\mathbb{N}}$.}$ (3.48)
Testing (2.13b) and (2.13c) written for $({\boldsymbol{v}}^{K_{n}},\varphi^{K_{n}},\mu^{K_{n}},\psi^{K_{n}},\theta^{K_{n}})$ by the constant functions $\frac{1}{|\Omega|}$ and $\frac{1}{|\Gamma|}$, respectively, we infer
${\langle\varphi^{K_{n}}(t)\rangle}_{\Omega}={\langle\varphi_{0}\rangle}_{\Omega}\quad\text{and}\quad{\langle\psi^{K_{n}}(t)\rangle}_{\Gamma}={\langle\psi_{0}\rangle}_{\Gamma}$
for all $t\in[0,T]$ and all $n\in{\mathbb{N}}$. Hence, proceeding analogously as in the proof of Theorem 2.4 (Subsection 3.1.2, First and Second estimates), we derive the uniform bound
$\left\|{\boldsymbol{v}}^{K_{n}}\right\|_{L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})\cap L^{2}(0,T;\mathbf{H}_{\Gamma})}+\left\|\nabla\mu^{K_{n}}\right\|_{L^{2}(0,T;{\bf{H}})}+\left\|\nabla_{\Gamma}\theta^{K_{n}}\right\|_{L^{2}(0,T;\mathbf{H}_{\Gamma})}$
$\quad+\left\|\varphi^{K_{n}}\right\|_{H^{1}(0,T;{V^{*}})\cap L^{\infty}(0,T;V)}+\left\|\psi^{K_{n}}\right\|_{H^{1}(0,T;{{V}_{\Gamma}^{*}})\cap L^{\infty}(0,T;{V}_{\Gamma})}\leq C.$ (3.49)
We now test (2.13d) written for $({\boldsymbol{v}}^{K_{n}},\varphi^{K_{n}},\mu^{K_{n}},\psi^{K_{n}},\theta^{K_{n}})$ by $(\eta,0)$, where $\eta\in C^{\infty}_{c}(\Omega)$ is an arbitrary test function. Using (3.2) along with Hölder’s inequality, we infer that
$\left|\int_{\Omega}\mu^{K_{n}}\eta\right|\leq\left\|\varphi^{K_{n}}\right\|_{V}\left\|\eta\right\|_{V}+\left\|F^{\prime}(\varphi^{K_{n}})\right\|_{L^{\frac{6}{5}}(\Omega)}\left\|\eta\right\|_{L^{6}(\Omega)}\leq C\left\|\eta\right\|_{V}$ (3.50)
entailing that $\mu^{K_{n}}$ is bounded in $L^{\infty}(0,T;V^{*})$. Then, by means of the generalized Poincaré inequality (see [8, Section 8.16] or [33, Lemma A.1]), we infer that
$\left\|\mu^{K_{n}}\right\|_{H}\leq C\big{(}1+\left\|\nabla\mu^{K_{n}}\right\|_{\bf{H}}\big{)}.$
For more details see, e.g., [33, Proof of Lemma 4.5]. Hence, in combination with (3.50), we conclude
$\left\|\mu^{K_{n}}\right\|_{L^{2}(0,T;V)}\leq C.$ (3.51)
In order to derive an analogous estimate for $\theta^{K_{n}}$, we first choose $\eta\equiv 1$ and ${\eta_{\Gamma}}\equiv 0$ in (2.13d). Employing (3.2), we obtain
$\left|\frac{1}{{K_{n}}}\int_{\Gamma}(\psi^{K_{n}}-\varphi^{K_{n}})\right|\leq\left\|\mu^{K_{n}}\right\|_{L^{1}(\Omega)}+\left\|F^{\prime}(\varphi^{K_{n}})\right\|_{L^{1}(\Omega)}\leq C\left\|\mu^{K_{n}}\right\|_{H}+C.$ (3.52)
Let us now take $\eta\equiv 0$ and ${\eta_{\Gamma}}\equiv 1$ in (2.13d). Using (3.2) and (3.52), we deduce
$\left|\int_{\Gamma}\theta^{K_{n}}\right|$ $\leq\left\|G^{\prime}(\psi^{K_{n}})\right\|_{L^{1}(\Gamma)}+\left|\frac{1}{{K_{n}}}\int_{\Gamma}(\psi^{K_{n}}-\varphi^{K_{n}})\right|\leq C\left\|\mu^{K_{n}}\right\|_{H}+C.$
Employing Poincaré’s inequality, we thus infer
$\left\|\theta^{K_{n}}\right\|_{{H}_{\Gamma}}\leq C\left\|\nabla_{\Gamma}\theta^{K_{n}}\right\|_{\mathbf{H}_{\Gamma}}+C\left\|\mu^{K_{n}}\right\|_{H}+C.$
Squaring and integrating this estimate with respect to time over $[0,T]$, we eventually conclude
$\left\|\theta^{K_{n}}\right\|_{L^{2}(0,T;{H}_{\Gamma})}\leq C.$ (3.53)
In summary, combining (3.2), (3.51) and (3.53), we have thus shown
$\left\|{\boldsymbol{v}}^{K_{n}}\right\|_{L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})\cap L^{2}(0,T;\mathbf{H}_{\Gamma})}+\left\|\mu^{K_{n}}\right\|_{L^{2}(0,T;V)}+\left\|\theta^{K_{n}}\right\|_{L^{2}(0,T;{V}_{\Gamma})}$
$\quad+\left\|\varphi^{K_{n}}\right\|_{H^{1}(0,T;{V^{*}})\cap L^{\infty}(0,T;V)}+\left\|\psi^{K_{n}}\right\|_{H^{1}(0,T;{{V}_{\Gamma}^{*}})\cap L^{\infty}(0,T;{V}_{\Gamma})}\leq C.$ (3.54)
As in Subsection 3.1.4, we deduce the existence of functions $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ such that the convergences (2.17a)–(2.17f) hold along a non-relabeled subsequence. Moreover, the estimate (3.48) directly implies (2.17g) and thus, all convergences in (2.5) are established. In particular, due to the trace theorem, we also have
$\varphi^{K_{n}}|_{\Gamma}-\psi^{K_{n}}\;\to\;\varphi^{0}|_{\Gamma}-\psi^{0}\quad\text{strongly in $C^{0}([0,T];{H}_{\Gamma})$}.$ (3.55)
In combination with (3.48), this proves that $\varphi^{0}|_{\Gamma}=\psi^{0}$ a.e. on $\Sigma$ due to uniqueness of the limit. Proceeding further as in Subsection 3.1.4, we eventually show that the quintuplet $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman system (1) in the sense of Definition 2.3.
It remains to verify the additional regularity assertions. Without loss of generality, we merely consider the case $d=3$. The case $d=2$ can be handled analogously but is even easier as the Sobolev embeddings in two dimensions are better. We infer from (2.13d) written for the solution $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ and $K=0$ that there exists a null set $\mathcal{N}\subset[0,T]$ such that
$\int_{\Omega}\nabla\varphi^{0}(t)\cdot\nabla\eta+\int_{\Gamma}\nabla_{\Gamma}\psi^{0}(t)\cdot\nabla_{\Gamma}{\eta_{\Gamma}}$
$\quad=\int_{\Omega}\Big{(}\mu^{0}(t)-F^{\prime}\big{(}\varphi^{0}(t)\big{)}\Big{)}\,\eta+\int_{\Gamma}\Big{(}\theta^{0}(t)-G^{\prime}\big{(}\psi^{0}(t)\big{)}\Big{)}\,{\eta_{\Gamma}}$ (3.56)
for all $t\in[0,T]\setminus\mathcal{N}$ and all test functions $(\eta,{\eta_{\Gamma}})\in{\cal V}_{0}$.
Let now $t\in[0,T]\setminus\mathcal{N}$ be arbitrary. We infer from (3.2) that the pair $\big{(}\varphi^{0}(t),\psi^{0}(t)\big{)}$ is a weak solution of the bulk-surface elliptic problem
$-\Delta\varphi^{0}(t)$ $=f(t)$ in $\quad\text{in $\Omega$},$ (3.57a)
$-\Delta_{\Gamma}\psi^{0}(t)+\partial_{\bf n}\varphi^{0}(t)$ $=g(t)$ on $\quad\text{on $\Gamma$},$ (3.57b)
$\varphi^{0}(t)|_{\Gamma}$ $=\psi^{0}(t)$ on $\quad\text{on $\Gamma$},$ (3.57c)
where
$f(t):=\mu^{0}(t)-F^{\prime}\big{(}\varphi^{0}(t)\big{)}\quad\text{and}\quad g(t):=\theta^{0}(t)-G^{\prime}\big{(}\psi^{0}(t)\big{)}.$
Let us first consider the case $\ell=2$. As we assumed that the growth conditions in (R1) are fulfilled with $p\leq 4$, we have
$\left\|F^{\prime}\big{(}\varphi^{0}(t)\big{)}\right\|_{H}$ $\leq C+C\left\|\varphi^{0}\right\|_{L^{6}(\Omega)}^{3}\leq C,$ (3.58)
$\left\|G^{\prime}\big{(}\psi^{0}(t)\big{)}\right\|_{{H}_{\Gamma}}$ $\leq C+C\left\|\psi^{0}\right\|_{L^{2(q-1)}(\Gamma)}^{q-1}\leq C.$ (3.59)
Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we find that $\big{(}\varphi^{0}(t),\psi^{0}(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
$\left\|\varphi^{0}(t)\right\|_{H^{2}(\Omega)}^{2}+\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}^{2}\leq C\left\|f^{0}(t)\right\|_{H}^{2}+C\left\|g^{0}(t)\right\|_{{H}_{\Gamma}}^{2}$
$\quad\leq C+C\left\|\mu^{0}(t)\right\|_{H}^{2}+C\left\|\theta^{0}(t)\right\|_{{H}_{\Gamma}}^{2}.$
Since $\mu^{0}\in L^{2}(0,T;H)$ and $\theta^{0}\in L^{2}(0,T;{H}_{\Gamma})$, this proves (2.18a).
We now consider the case $\ell=3$. Recalling that the growth conditions in (R1) are fulfilled with $p\leq 4$, we use (3.2) to derive the estimates
$\left\|F^{\prime\prime}\big{(}\varphi^{0}(t)\big{)}\nabla\varphi^{0}(t)\right\|_{\bf H}$ $\leq C\left\|\nabla\varphi^{0}(t)\right\|_{{\bf H}}+C\left\||\varphi^{0}(t)|^{2}\,\nabla\varphi^{0}(t)\right\|_{{\bf H}}$
$\leq C+C\left\|\varphi^{0}(t)\right\|_{L^{6}(\Omega)}^{2}\,\left\|\nabla\varphi^{0}(t)\right\|_{L^{6}(\Omega)}$
$\leq C+C\left\|\varphi^{0}(t)\right\|_{H^{2}(\Omega)}$
and
$\left\|G^{\prime\prime}\big{(}\psi^{0}(t)\big{)}\nabla_{\Gamma}\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}$ $\leq C\left\|\nabla\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}+C\left\||\psi^{0}(t)|^{q-2}\,\nabla\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}$
$\leq C+C\left\|\psi^{0}(t)\right\|_{L^{{2(q-2)}}}^{q-2}\left\|\nabla\psi^{0}(t)\right\|_{L^{4}(\Gamma)}$
$\leq C+C\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}.$
In combination with (3.58) and (3.59), these estimates directly imply
$\left\|F^{\prime}\big{(}\varphi^{0}(t)\big{)}\right\|_{V}$ $\leq C+C\left\|\varphi^{0}\right\|_{H^{2}(\Omega)},$
$\left\|G^{\prime}\big{(}\psi^{0}(t)\big{)}\right\|_{{V}_{\Gamma}}$ $\leq C+C\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}.$
Now, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we infer
$\left\|\varphi^{0}(t)\right\|_{H^{3}(\Omega)}^{2}+\left\|\psi^{0}(t)\right\|_{H^{3}(\Gamma)}^{2}\leq C\left\|f^{0}(t)\right\|_{V}^{2}+C\left\|g^{0}(t)\right\|_{{V}_{\Gamma}}^{2}$
$\quad\leq C+C\left\|\mu^{0}(t)\right\|_{V}^{2}+C\left\|\theta^{0}(t)\right\|_{{V}_{\Gamma}}^{2}+C\left\|\varphi^{0}(t)\right\|_{H^{2}(\Omega)}^{2}+C\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}^{2}.$
Recalling $\mu^{0}\in L^{2}(0,T;V)$, $\theta^{0}\in L^{2}(0,T;{V}_{\Gamma})$ and that (2.18a) with $\ell=2$ is already verified, this proves (2.18b) in the case $\ell=3$. By means of Proposition A.1(b), we directly infer $(\varphi^{0},\psi^{0})\in C^{0}([0,T];{\cal V}_{0})$. Moreover, via interpolation between $L^{\infty}(0,T;{\cal V}_{0})$ and $L^{2}(0,T;H^{3}(\Omega)\times H^{3}(\Gamma))$ (cf. Lemma 2.2), we further get
$(\varphi^{0},\psi^{0})\in L^{4}(0,T;H^{2}(\Omega)\times H^{2}(\Gamma)).$
This means that (2.19) is established. Eventually, a simple comparison argument based on (2.13d) yields (2.20).
This means that all assertions are verified and thus, the proof is complete. ∎
3.3   Uniqueness of the weak solution for regular potentials
In this subsection, we are going to prove Theorem 2.7 for regular potentials and $K\geq 0$. To prove the theorem, we use some ideas devised in [40].
Proof of Theorem 2.7.
In this proof, the the letter $C$ will denote generic positive constants that may depend on $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3). Such constants may also change their value from line to line. We first introduce the following notation for the differences of the solution components:
${\boldsymbol{v}}:={\boldsymbol{v}}_{1}-{\boldsymbol{v}}_{2},\quad\varphi:=\varphi_{1}-\varphi_{2},\quad\mu:=\mu_{1}-\mu_{2},\quad\psi:=\psi_{1}-\psi_{2},\quad\theta:=\theta_{1}-\theta_{2}.$
Recall that $M,M_{\Gamma}$ and $\nu$ are assumed to be constant. We thus infer that the quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ satisfies the variational formulation
$2\nu\int_{\Omega}\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}(\lambda(\varphi_{1})-\lambda(\varphi_{2})){\boldsymbol{v}}_{1}\cdot{\boldsymbol{w}}$
$\qquad+\int_{\Omega}\lambda(\varphi_{2}){\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}(\gamma(\psi_{1})-\gamma(\psi_{2})){\boldsymbol{v}}_{1}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma(\psi_{2}){\boldsymbol{v}}\cdot{\boldsymbol{w}}$
$\quad=-\int_{\Omega}(\varphi\nabla\mu_{1}+\varphi_{2}\nabla\mu)\cdot{\boldsymbol{w}}-\int_{\Gamma}(\psi\nabla_{\Gamma}\theta_{1}+\psi_{2}\nabla_{\Gamma}\theta)\cdot{\boldsymbol{w}},$ (3.60a)
$\mathopen{\langle}\partial_{t}\varphi,\zeta\mathclose{\rangle}_{V}-\int_{\Omega}(\varphi{\boldsymbol{v}}_{1}+\varphi_{2}{\boldsymbol{v}})\cdot\nabla\zeta+{M}\int_{\Omega}\nabla\mu\cdot\nabla\zeta=0,$ (3.60b)
$\mathopen{\langle}\partial_{t}\psi,{\zeta_{\Gamma}}\mathclose{\rangle}_{{V}_{\Gamma}}-\int_{\Gamma}(\psi{\boldsymbol{v}}_{1}+\psi_{2}{\boldsymbol{v}})\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}+{M_{\Gamma}}\int_{\Gamma}\nabla_{\Gamma}\theta\cdot\nabla_{\Gamma}{\zeta_{\Gamma}}=0$ (3.60c)
almost everywhere in $(0,T)$ for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, and ${\zeta_{\Gamma}}\in{V}_{\Gamma}$, and the equations
$\mu$ $=-\Delta\varphi+F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\quad$ in $Q$, (3.61)
$\theta$ $=-\Delta_{\Gamma}\psi+G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})+\partial_{\bf n}\varphi\quad$ on $\Sigma$ (3.62)
are fulfilled in the strong sense due the higher regularities established in Theorem 2.4 and Theorem 2.5.
We now test (3.60a) by ${\boldsymbol{v}}$, (3.60b) by $\varphi+\mu$, (3.60c) by $\psi+\theta$, and add the resulting equations. After some cancellations and rearrangements, we obtain
$2\nu\left\|\mathbf{D}{\boldsymbol{v}}\right\|^{2}_{\mathbb{H}}+\int_{\Omega}\lambda(\varphi_{2})|{{\boldsymbol{v}}}|^{2}+\int_{\Gamma}\gamma(\psi_{2})|{{\boldsymbol{v}}}|^{2}+\mathopen{\langle}\partial_{t}\varphi,\varphi+\mu\mathclose{\rangle}_{V}$
$\qquad+{M}\left\|\nabla\mu\right\|^{2}_{{\bf H}}+\mathopen{\langle}\partial_{t}\psi,\psi+\theta\mathclose{\rangle}_{{V}_{\Gamma}}+{M_{\Gamma}}\left\|\nabla_{\Gamma}\theta\right\|^{2}_{{\bf H}_{\Gamma}}$
$\quad=-\int_{\Omega}(\lambda(\varphi_{1})-\lambda(\varphi_{2})){\boldsymbol{v}}_{1}\cdot{\boldsymbol{v}}-\int_{\Gamma}(\gamma(\psi_{1})-\gamma(\psi_{2})){\boldsymbol{v}}_{1}\cdot{\boldsymbol{v}}$
$\qquad-\int_{\Omega}\varphi\nabla\mu_{1}\cdot{\boldsymbol{v}}-\int_{\Gamma}\psi\nabla_{\Gamma}\theta_{1}\cdot{\boldsymbol{v}}$
$\qquad+\int_{\Omega}(\varphi{\boldsymbol{v}}_{1}+\varphi_{2}{\boldsymbol{v}})\cdot\nabla\varphi+\int_{\Omega}\varphi{\boldsymbol{v}}_{1}\cdot\nabla\mu$
$\qquad+\int_{\Gamma}(\psi{\boldsymbol{v}}_{1}+\psi_{2}{\boldsymbol{v}})\cdot\nabla_{\Gamma}\psi+\int_{\Gamma}\psi{\boldsymbol{v}}_{1}\cdot\nabla_{\Gamma}\theta$
$\qquad-{M}\int_{\Omega}\nabla\mu\cdot\nabla\varphi-{M_{\Gamma}}\int_{\Gamma}\nabla_{\Gamma}\theta\cdot\nabla_{\Gamma}\psi=:\sum_{i=1}^{10}I_{i}.$ (3.63)
We point out that, as a consequence of Theorem 2.4 and Theorem 2.5, it holds that $(\varphi_{i},\psi_{i})\in L^{2}(0,T;(H^{3}(\Omega)\times H^{3}(\Gamma))\cap{\cal V}_{K})$, $i=1,2$. Next, by using (3.61) and (3.62), along with the chain rule formula in Proposition A.1, we observe that the duality terms on the left-hand side can be reformulated as
$\mathopen{\langle}\partial_{t}\varphi,\varphi+\mu\mathclose{\rangle}_{V}+\mathopen{\langle}\partial_{t}\psi,\psi+\theta\mathclose{\rangle}_{{V}_{\Gamma}}$
$\quad=\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\Big{(}{\left\|\varphi\right\|^{2}_{H}+\left\|\psi\right\|^{2}_{{H}_{\Gamma}}}\Big{)}+\mathopen{\langle}(\partial_{t}\varphi,\partial_{t}\psi),\big{(}-\Delta\varphi,-\Delta_{\Gamma}\psi+\partial_{\bf n}\varphi\big{)}\mathclose{\rangle}_{{\cal V}}$
$\qquad+\mathopen{\langle}\partial_{t}\varphi,F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\mathclose{\rangle}_{V}+\mathopen{\langle}\partial_{t}\psi,G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\mathclose{\rangle}_{{V}_{\Gamma}}$
$\quad=\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\Big{(}{\left\|\varphi\right\|^{2}_{V}+\left\|\psi\right\|^{2}_{{V}_{\Gamma}}}+\sigma(K)\left\|\psi-\varphi\right\|^{2}_{{H}_{\Gamma}}\Big{)}$
$\qquad+\mathopen{\langle}\partial_{t}\varphi,F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\mathclose{\rangle}_{V}+\mathopen{\langle}\partial_{t}\psi,G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\mathclose{\rangle}_{{V}_{\Gamma}}.$ (3.64)
Using (A2) and (A3) as well as (3.64), we deduce from (3.63) that
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\Big{(}\left\|\varphi\right\|_{V}^{2}+\left\|\psi\right\|_{{V}_{\Gamma}}^{2}+\sigma(K)\left\|\psi-\varphi\right\|^{2}_{{H}_{\Gamma}}\Big{)}$
$\quad+2\nu\left\|\mathbf{D}{\boldsymbol{v}}\right\|^{2}_{\mathbb{H}}+\lambda_{1}\left\|{\boldsymbol{v}}\right\|_{{\bf H}}^{2}+\gamma_{1}\left\|{\boldsymbol{v}}\right\|_{{\bf H}_{\Gamma}}^{2}+{M}\left\|\nabla\mu\right\|^{2}_{{\bf H}}+{M_{\Gamma}}\left\|\nabla_{\Gamma}\theta\right\|^{2}_{{\bf H}_{\Gamma}}$
$\qquad\leq-\mathopen{\langle}\partial_{t}\varphi,F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\mathclose{\rangle}_{V}-\mathopen{\langle}\partial_{t}\psi,G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\mathclose{\rangle}_{{V}_{\Gamma}}+\sum_{i=1}^{10}I_{i}.$
We now intend to control the terms $I_{i}$, $i=1,...,10$, by means of Hölder’s inequality, Young’s inequality, the Lipschitz continuity of $\lambda$ and $\gamma$, and integration by parts along with Sobolev’s embeddings and the trace theorem. For a positive $\delta$ yet to be chosen, we derive the following estimates:
${I_{1}}$ $\leq C\left\|\varphi\right\|_{{L^{4}(\Omega)}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Omega)}}\left\|{\boldsymbol{v}}\right\|_{{{\bf H}}}\leq\delta\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf H}}+C_{\delta}\left\|{\boldsymbol{v}}_{1}\right\|^{2}_{{\bf V}}\left\|\varphi\right\|_{V}^{2},$
${I_{2}}$ $\leq C\left\|\psi\right\|_{{L^{4}(\Gamma)}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Gamma)}}\left\|{\boldsymbol{v}}\right\|_{{\mathbf{H}_{\Gamma}}}\leq\delta\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf{V}}}+C_{\delta}\left\|{\boldsymbol{v}}_{1}\right\|^{2}_{{\bf{V}}}\left\|\psi\right\|^{2}_{{V}_{\Gamma}},$
${I_{3}+I_{4}}$ $=\int_{\Omega}\mu_{1}\nabla\varphi\cdot{\boldsymbol{v}}-\int_{\Gamma}\psi\nabla_{\Gamma}\theta_{1}\cdot{\boldsymbol{v}}$
${\leq\left\|\mu_{1}\right\|_{{L^{4}(\Omega)}}\left\|\nabla\varphi\right\|_{{{\bf{H}}}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{L}}^{4}(\Omega)}}+\left\|\psi\right\|_{{L^{4}(\Gamma)}}\left\|\nabla_{\Gamma}\theta_{1}\right\|_{{\mathbf{H}_{\Gamma}}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{L}}^{4}(\Gamma)}}}$
${\leq{C}\left\|\mu_{1}\right\|_{V}\left\|\varphi\right\|_{V}\left\|{\boldsymbol{v}}\right\|_{{\bf V}}+C\left\|\theta_{1}\right\|_{{V}_{\Gamma}}\left\|\psi\right\|_{{V}_{\Gamma}}\left\|{\boldsymbol{v}}\right\|_{\bf V}}$
${\leq 2\delta\left\|{\boldsymbol{v}}\right\|^{2}_{\bf V}+C_{\delta}\left\|\mu_{1}\right\|^{2}_{V}\left\|\varphi\right\|^{2}_{V}+C_{\delta}\left\|\theta_{1}\right\|_{{V}_{\Gamma}}^{2}\left\|\psi\right\|_{{V}_{\Gamma}}^{2},}$
${I_{5}+I_{6}}$ $\leq(\left\|\varphi\right\|_{{L^{4}(\Omega)}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Omega)}}+\left\|\varphi_{2}\right\|_{{L^{4}(\Omega)}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{L}}^{4}(\Omega)}})\left\|\nabla\varphi\right\|_{{{\bf{H}}}}$
$\quad+\left\|\varphi\right\|_{L^{4}(\Omega)}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Omega)}}\left\|\nabla\mu\right\|_{{\bf H}}$
$\leq\frac{M}{4}\left\|\nabla\mu\right\|_{{\bf H}}^{2}+\delta\left\|{\boldsymbol{v}}\right\|_{{\bf V}}^{2}+(C\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf{V}}}+C\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf{V}}}^{2}+C_{\delta}\left\|\varphi_{2}\right\|_{V}^{2})\left\|\varphi\right\|_{V}^{2},$
${I_{7}+I_{8}}$ $\leq{(\left\|\psi\right\|_{{L^{4}(\Gamma)}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Gamma)}}+\left\|\psi_{2}\right\|_{{L^{4}(\Gamma)}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{L}}^{4}(\Gamma)}})\left\|\nabla_{\Gamma}\psi\right\|_{{\mathbf{H}_{\Gamma}}}}$
$\quad+\left\|\psi\right\|_{L^{4}(\Gamma)}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Gamma)}}\left\|\nabla_{\Gamma}\theta\right\|_{\mathbf{H}_{\Gamma}}$
${\leq\frac{M_{\Gamma}}{4}\left\|\nabla_{\Gamma}\theta\right\|_{\mathbf{H}_{\Gamma}}^{2}+\delta\left\|{\boldsymbol{v}}\right\|_{{\bf V}}^{2}+(C\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf{V}}}+C\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf{V}}}^{2}+C_{\delta}\left\|\psi_{2}\right\|_{{V}_{\Gamma}}^{2})\left\|\psi\right\|_{{V}_{\Gamma}}^{2},}$
${I_{9}+I_{10}}$ $\leq{\frac{{M}}{4}\left\|\nabla\mu\right\|^{2}_{{\bf H}}+\frac{{M_{\Gamma}}}{4}\left\|\nabla_{\Gamma}\theta\right\|^{2}_{{\bf H}_{\Gamma}}+C\left\|\nabla\varphi\right\|^{2}_{{\bf H}}+C\left\|\nabla_{\Gamma}\psi\right\|^{2}_{{\bf H}_{\Gamma}}.}$
Furthermore, the terms in the last line of (3.64) can be estimated by
$\big{|}\mathopen{\langle}\partial_{t}\varphi,F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\mathclose{\rangle}_{V}\big{|}+\big{|}\mathopen{\langle}\partial_{t}\psi,G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\mathclose{\rangle}_{{V}_{\Gamma}}\big{|}$
$\quad\leq\left\|\partial_{t}\varphi\right\|_{V^{*}}\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{V}+\left\|\partial_{t}\psi\right\|_{{V}_{\Gamma}^{*}}\left\|G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\right\|_{{V}_{\Gamma}}.$ (3.65)
By means of a comparison argument in (3.60b), we obtain
$\left\|\partial_{t}\varphi\right\|_{V^{*}}$ $=\sup_{\left\|\zeta\right\|_{{V}}\leq 1}|\mathopen{\langle}\partial_{t}\varphi,\zeta\mathclose{\rangle}_{V}|$
$\leq C(\left\|\varphi\right\|_{{L^{4}(\Omega)}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{L}}^{4}(\Omega)}}+\left\|\varphi_{2}\right\|_{{L^{4}(\Omega)}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{L}}^{4}(\Omega)}}+\left\|\nabla\mu\right\|_{{\bf H}})$
$\leq C(\left\|\varphi\right\|_{V}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{V}}}}+\left\|\varphi_{2}\right\|_{{V}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{V}}}}+\left\|\nabla\mu\right\|_{{\bf H}}).$
Similarly, using (3.60c), we derive the estimate
$\left\|\partial_{t}\psi\right\|_{{V}_{\Gamma}^{*}}=\sup_{\left\|\zeta_{\Gamma}\right\|_{{{V}_{\Gamma}}}\leq 1}|\mathopen{\langle}\partial_{t}\psi,\zeta_{\Gamma}\mathclose{\rangle}_{{V}_{\Gamma}}|\leq C(\left\|\psi\right\|_{{V}_{\Gamma}}\left\|{\boldsymbol{v}}_{1}\right\|_{{{\bf{V}}}}+\left\|\psi_{2}\right\|_{{{V}_{\Gamma}}}\left\|{\boldsymbol{v}}\right\|_{{{\bf{V}}}}+\left\|\nabla_{\Gamma}\theta\right\|_{{\bf H}_{\Gamma}}).$
By employing equation (3.60b) as well as (Renumi), we can bound the norm on the right-hand side of (3.65) as follows:
$\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{V}^{2}=\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{H}^{2}+\left\|F^{\prime\prime}(\varphi_{1})\nabla\varphi_{1}-F^{\prime\prime}(\varphi_{2})\nabla\varphi_{2}\right\|_{{\bf{H}}}^{2}$
$\quad=\int_{\Omega}|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})|^{2}+\int_{\Omega}|F^{\prime\prime}(\varphi_{1})\nabla\varphi|^{2}+\int_{\Omega}|F^{\prime\prime}(\varphi_{1})-F^{\prime\prime}(\varphi_{2})|^{2}|\nabla\varphi_{2}|^{2}$
$\quad\leq\int_{\Omega}\Big{|}\int_{0}^{1}F^{\prime\prime}\big{(}s\varphi_{1}+(1-s)\varphi_{2}\big{)}{\rm ds}\Big{|}^{2}\varphi^{2}\;+\;C(\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}+1)\left\|\nabla\varphi\right\|^{2}_{{\bf H}}$
$\qquad+\int_{\Omega}\Big{|}\int_{0}^{1}F^{(3)}\big{(}s\varphi_{1}+(1-s)\varphi_{2}\big{)}{\rm ds}\Big{|}^{2}\varphi^{2}\,|\nabla\varphi_{2}|^{2}$
$\quad\leq C(\left\|\varphi_{1}\right\|^{2(p-2)}_{{L^{3(p-2)}(\Omega)}}+\left\|\varphi_{2}\right\|^{2(p-2)}_{{L^{3(p-2)}(\Omega)}}+1)\left\|\varphi\right\|_{V}^{2}+C(\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}+1)\left\|\varphi\right\|^{2}_{V}$
$\qquad+C(\left\|\varphi_{1}\right\|^{2(p-3)}_{{L^{12(p-3)}(\Omega)}}+\left\|\varphi_{2}\right\|^{2(p-3)}_{{L^{12(p-3)}(\Omega)}}+1)\left\|\varphi_{2}\right\|_{W^{1,4}(\Omega)}^{2}\left\|\varphi\right\|_{V}^{2}.$ (3.66)
We now recall the restrictions on $p$ and $q$ demanded in (2.21). In particular, we have $p\leq 6$ if $d=3$. In the case $d=2$ we assume, without loss of generality, that $p\geq 5$. Using Agmon’s inequality as well as interpolation between Sobolev spaces (see Lemma 2.2), we derive the estimates
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{s}(\Omega)}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{2p-8}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ for $\text{for $d=2$},$
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{(p-2)}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ for $\text{for $d=3$},$
$\left\|\varphi_{2}\right\|_{W^{1,4}(\Omega)}^{2}\leq C\left\|\varphi_{2}\right\|_{H^{1}(\Omega)}^{\frac{1}{2}}\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}\leq C\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}$ for $\text{for $d=2,3$},$
where, in the first inequality, $s=\frac{2p}{2(p-2)}\in(1,2)$. We thus infer from (3.3) that
$\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{V}^{2}\leq C\Lambda\left\|\varphi\right\|_{V}^{2}$
with a the time-dependent function $\Lambda$ that is given by
$\Lambda$ $:=(1+\left\|\varphi_{1}\right\|^{4}_{H^{2}{(\Omega)}})+\big{(}1+\left\|\varphi_{2}\right\|^{\frac{3}{2}}_{H^{2}{(\Omega)}}\big{)}\sum_{i=1,2}\big{(}1+\left\|\varphi_{i}\right\|^{2(p-2)}_{{L^{3(p-2)}(\Omega)}}+\left\|\varphi_{i}\right\|^{2(p-3)}_{{L^{12(p-3)}(\Omega)}}\big{)}.$
From (2.15a) and (2.19), we know that $\varphi_{2}\in L^{4}(0,T;H^{2}(\Omega))$.
In the case $d=2$, we simply have
$\varphi_{i}\in L^{\infty}(0,T;L^{3(p-2)}(\Omega))\cap L^{\infty}(0,T;L^{12(p-3)}(\Omega)),\quad i=1,2,$
due to the Sobolev embedding $H^{1}(\Omega){\hookrightarrow}L^{r}(\Omega)$ for all $r\in(1,\infty)$.
In the case $d=3$, we use interpolation between Sobolev spaces (Lemma 2.2) to derive the embedding
$L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{3}(\Omega)){\hookrightarrow}L^{\rho}(0,T;L^{\frac{6\rho}{\rho-8}}(\Omega))\quad\text{for any $\rho>8$.}$ (3.67)
Since $\varphi_{i}\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{3}(\Omega))$, $i=1,2$, we infer
$\varphi_{i}\in L^{16}(0,T;L^{12}(\Omega))\cap L^{\frac{48}{5}}(0,T;L^{36}(\Omega)),\quad i=1,2,$
by choosing $\rho=16$ and $\rho=\frac{48}{5}$ in (3.67), respectively.
In summary, by means of Hölder’s inequality, we conclude
$t\mapsto{\Lambda(t)}\in L^{1}(0,T)\quad\text{for $d=2,3$.}$
Arguing in a similar fashion, and recalling (2.23) as well as the regularity in Theorem 2.4, we find that
$\left\|G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\right\|_{{V}_{\Gamma}}^{2}$
$\quad\leq(\left\|\psi_{1}\right\|^{2(q-2)}_{{V}_{\Gamma}}+\left\|\psi_{2}\right\|^{2(q-2)}_{{V}_{\Gamma}}+1)\left\|\psi\right\|_{{V}_{\Gamma}}^{2}+C(\left\|\psi_{1}\right\|_{L^{\infty}(\Gamma)}^{2(q-2)}+1)\left\|\psi\right\|_{{V}_{\Gamma}}^{2}$
$\qquad+C\big{(}1+\left\|\psi_{1}\right\|_{{V}_{\Gamma}}^{{2(q-3)}}+\left\|\psi_{2}\right\|_{{V}_{\Gamma}}^{2(q-3)}\big{)}\left\|\psi_{2}\right\|_{H^{2}(\Gamma)}^{2}\left\|\psi\right\|^{2}_{{V}_{\Gamma}}$
$\quad\leq C(\left\|\psi_{1}\right\|_{L^{\infty}(\Gamma)}^{2(q-2)}+\left\|\psi_{2}\right\|_{H^{2}(\Gamma)}^{2}+1)\left\|\psi\right\|_{{V}_{\Gamma}}^{2}.$
In view of (2.21), we assume, without loss of generality, that $q\geq 5$. Recalling that the boundary $\Gamma$ is a $(d-1)$-dimensional submanifold of ${\mathbb{R}}^{d}$ with $d\in\{2,3\}$, we have $H^{s}(\Gamma){\hookrightarrow}L^{\infty}(\Gamma)$ for every $s>1$. Hence, via interpolation between Sobolev spaces (see Lemma 2.2) we obtain the estimate
$\left\|\psi_{1}\right\|_{L^{\infty}(\Gamma)}^{2(q-2)}\leq C\left\|\psi_{1}\right\|_{H^{s}(\Gamma)}^{2(q-2)}\leq C\left\|\psi_{1}\right\|_{{V}_{\Gamma}}^{2q-8}\left\|\psi_{1}\right\|_{H^{2}(\Gamma)}^{4}\leq C\left\|\psi_{1}\right\|_{H^{2}(\Gamma)}^{4},$
where $s=\frac{2q}{2(q-2)}\in(1,2)$. We thus conclude that
$\left\|G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\right\|_{{V}_{\Gamma}}^{2}\leq C\Theta\left\|\psi\right\|_{{V}_{\Gamma}}^{2}$
with a time-dependent function $\Theta$ that is given by
$t\mapsto\Theta(t):=C(1+\left\|\psi_{1}(t)\right\|^{4}_{H^{2}(\Gamma)}+\left\|\psi_{2}(t)\right\|^{2}_{H^{2}(\Gamma)})\in L^{1}(0,T).$
Therefore, upon collecting the above computations, the integral in (3.65) can be estimated with the help of Young’ inequality as
$-\mathopen{\langle}\partial_{t}\varphi,F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\mathclose{\rangle}_{V}-\mathopen{\langle}\partial_{t}\psi,G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\mathclose{\rangle}_{{V}_{\Gamma}}$
$\quad\leq C\left\|\partial_{t}\varphi\right\|_{V^{*}}\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{V}+C\left\|\partial_{t}\psi\right\|_{{V}_{\Gamma}^{*}}\left\|G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\right\|_{{V}_{\Gamma}}$
$\quad\leq\delta(\left\|\partial_{t}\varphi\right\|_{V^{*}}^{2}+\left\|\partial_{t}\psi\right\|_{{V}_{\Gamma}^{*}}^{2})+C_{\delta}(\left\|F^{\prime}(\varphi_{1})-F^{\prime}(\varphi_{2})\right\|_{V}^{2}+\left\|G^{\prime}(\psi_{1})-G^{\prime}(\psi_{2})\right\|_{{V}_{\Gamma}}^{2})$
$\quad\leq\delta C(\left\|\nabla\mu\right\|^{2}_{{\bf H}}+\left\|\nabla_{\Gamma}\theta\right\|^{2}_{\mathbf{H}_{\Gamma}})+\delta C\left\|\varphi_{2}\right\|_{L^{\infty}(0,T;V)}^{2}\left\|{\boldsymbol{v}}\right\|_{{\bf V}}^{2}+C_{\delta}(\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf V}}^{2}+\Lambda)\left\|\varphi\right\|_{V}^{2}$
$\qquad+\delta C\left\|\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}^{2}\left\|{\boldsymbol{v}}\right\|_{{\bf V}}^{2}+C_{\delta}(\left\|{\boldsymbol{v}}_{1}\right\|_{{\bf V}}^{2}+\Theta)\left\|\psi\right\|_{{V}_{\Gamma}}^{2}$
for a constant $\delta>0$ yet to be chosen. Finally, we adjust $\delta\in(0,1)$ in such a way that
$\delta\max\Big{\{}4,C,C\left\|\varphi_{2}\right\|_{L^{\infty}(0,T;V)}^{2},C\left\|\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}^{2}\Big{\}}\leq\frac{1}{4}\min\Big{\{}{{M}},{{M_{\Gamma}}},{C_{P}(\nu,\gamma_{1})}\Big{\}},$
where $C_{P}(\nu,\gamma_{1})$ is a Poincaré constant such that $2\nu\left\|\mathbf{D}{\boldsymbol{v}}\right\|^{2}_{\mathbb{H}}+\gamma_{1}\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf H}_{\Gamma}}\geq C_{P}(\nu,\gamma_{1})\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf{V}}}$. Thus, we integrate over time and employ Gronwall’s lemma to deduce that
$\left\|{\boldsymbol{v}}_{1}-{\boldsymbol{v}}_{2}\right\|_{L^{2}(0,T;{\bf{V}})}+\left\|\varphi_{1}-\varphi_{2}\right\|_{L^{\infty}(0,T;V)}+\left\|\nabla\mu_{1}-\nabla\mu_{2}\right\|_{L^{2}(0,T;{\bf{H}})}$
$\qquad+\left\|\psi_{1}-\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}+\left\|\nabla_{\Gamma}\theta_{1}-\nabla_{\Gamma}\theta_{2}\right\|_{L^{2}(0,T;\mathbf{H}_{\Gamma})}$
$\quad\leq C(\left\|\varphi_{0,1}-\varphi_{0,2}\right\|_{V}+\left\|\psi_{0,1}-\psi_{0,2}\right\|_{{V}_{\Gamma}}).$
Finally, by a comparison argument in (3.61) and (3.62), we infer that $(\mu,\theta)$ is bounded in $L^{2}(0,T;{\cal H})$ by the same right-hand side as the above inequality. This leads to (2.24) and thus, the proof is complete. ∎
4   Analysis of the Cahn–Hilliard–Brinkman system with singular potentials
We are now dealing with the proof of the existence of weak solutions for singular potentials. Our strategy is to approximate the convex parts of the singular potentials $F$ and $G$ satisfying (S1) and (S2) by means of a Yosida regularization. In this way, the approximate potentials are regular and exhibit quadratic growth and we can thus use Theorem 2.4 and Theorem 2.5 to obtain suitable approximate solutions. We then derive uniform estimates with respect to the approximation parameter, and eventually pass to the limit. In the forthcoming analysis, the splitting $F^{\prime}=\beta+\pi$ and $G^{\prime}={\beta_{\Gamma}}+\pi_{\Gamma}$ from (S1) will be adopted.
4.1   Yosida regularization
As mentioned, we rely on a Yosida regularization to smooth the singular parts of the potentials $F$ and $G$. For any $\varepsilon\in(0,1)$, we approximate the maximal monotone graphs $\beta$ and ${\beta_{\Gamma}}$ by
${\beta_{\varepsilon}}(r)$ $:=\frac{1}{\varepsilon}\Big{(}r-\big{(}I+\varepsilon\beta\big{)}^{-1}(r)\Big{)},\quad\beta_{\Gamma,\varepsilon}(r):=\frac{1}{\varepsilon}\Big{(}r-\big{(}I+\varepsilon{\beta_{\Gamma}}\big{)}^{-1}(r)\Big{)},\quad r\in{\mathbb{R}}.$
Then, the condition in (2.26) implies that
$\bigl{|}\beta_{\varepsilon}(r)\bigr{|}\leq\kappa_{1}\bigl{|}\beta_{\Gamma,\varepsilon}(r)\bigr{|}+\kappa_{2}\quad\text{for all $r\in\mathbb{R}$ and all $\varepsilon\in(0,1)$}$ (4.1)
(see, e.g., [15, Appendix]), where $\kappa_{1}$ and $\kappa_{2}$ are the constants introduced in (2.26). Next, we define $F_{\varepsilon}:=\widehat{\beta_{\varepsilon}}+\widehat{\pi},$ $G_{\varepsilon}:=\widehat{\beta}_{\Gamma,\varepsilon}+\widehat{\pi}_{\Gamma},$ where
$\widehat{\beta_{\varepsilon}}(r):=\int_{0}^{r}{\beta_{\varepsilon}}(s)\,{\rm ds},\quad\widehat{\beta}_{\Gamma,\varepsilon}(r):=\int_{0}^{r}\beta_{\Gamma,\varepsilon}(s)\,{\rm ds},\quad r\in{\mathbb{R}}.$
It is well-known that for every $r\in{\mathbb{R}}$,
$0\leq\widehat{\beta_{\varepsilon}}(r)\leq\widehat{\beta}(r)$ $\quad\forall\varepsilon\in(0,1),\quad$ $\widehat{\beta_{\varepsilon}}(r)$ $\nearrow\widehat{\beta}(r)$ monotonically as $\varepsilon\to 0,$ (4.2a)
$|{\beta_{\varepsilon}}(r)|\leq|\beta^{\circ}(r)|$ $\quad\forall\varepsilon\in(0,1),\quad$ ${\beta_{\varepsilon}}(r)$ $\to\beta^{\circ}(r)$ as $\varepsilon\to 0$. (4.2b)
Analogous properties hold for $\beta_{\Gamma,\varepsilon}$. Moreover, owing to the growth condition (2.25), $\widehat{\beta_{\varepsilon}}$ fulfills the following growth condition:
$&\hbox{For every $M>0$ there exist $C_{M}>0$ and $\varepsilon_{M}\in(0,1)$ such that}\\ &\widehat{\beta_{\varepsilon}}(r)\geq M\,r^{2}-C_{M}\quad\hbox{for every $r\in{\mathbb{R}}$ and every $\varepsilon\in(0,\varepsilon_{M})$}\,.\end{aligned}$ (4.3)
This property is checked in detail in the paper [20, beginning of Section 3]. Obviously, as a consequence, a similar condition holds for $\widehat{\beta}_{\Gamma,\varepsilon}$ since (4.1) entails that
$\widehat{\beta_{\varepsilon}}(r)\leq\kappa_{1}\widehat{\beta}_{\Gamma,\varepsilon}(r)+\kappa_{2}|r|\quad\text{for every $r\in{\mathbb{R}},\ \varepsilon\in(0,1)$},$ (4.4)
thanks to ${\beta_{\varepsilon}}(0)=\beta_{\Gamma,\varepsilon}(0)=0$ and since $\widehat{\beta_{\varepsilon}}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have the same sign. Due to their construction by the Yosida approximation, $\beta_{\varepsilon}$ and $\beta_{\Gamma,\varepsilon}$ are Lipschitz continuous and have at most linear growth. Hence, $\widehat{\beta}_{\varepsilon}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have at most quadratic growth. Moreover, (4.3) and (4.4) along with the (at most) quadratic growth of $\widehat{\pi}$ and ${\widehat{\pi}}_{\Gamma}$ (cf. (S1)), imply that both $F_{\varepsilon}$ and $G_{\varepsilon}$ are bounded from below by negative constants independent of $\varepsilon$. We can thus assume, without loss of generality, that $F_{\varepsilon}$ and $G_{\varepsilon}$ are nonnegative (otherwise, we add the modulus of their lower bounds, respectively). This entails that the approximate potentials $F_{\varepsilon}$ and $G_{\varepsilon}$ satisfy assumption (R1) with $p=q=2$.
Now, the approximating system we aim to solve consists of (2.28a)–(2.28d) with $\beta={\beta_{\varepsilon}}$ and ${\beta_{\Gamma}}=\beta_{\Gamma,\varepsilon}$. The regularity of the approximate potentials, in particular, implies that the inclusions $\xi_{\varepsilon}\in{\beta_{\varepsilon}}(\varphi_{\varepsilon})$ a.e. in $Q$ and $\xi_{\Gamma,\varepsilon}\in\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$ a.e. on $\Sigma$ turn into the identities $\xi_{\varepsilon}={\beta_{\varepsilon}}(\varphi_{\varepsilon})$ a.e. in $Q$and $\xi_{\Gamma,\varepsilon}=\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$ a.e. on $\Sigma$, respectively.
Therefore, as an immediate consequence of Theorem 2.4 and Theorem 2.5, we obtain the following existence result.
Corollary 4.1.
Let $K\geq 0$, suppose that (A1)(A3) hold, and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be arbitrary initial data. Then, for every $\varepsilon\in(0,1)$, the approximate problem described above admits at least a weak solution $({\boldsymbol{v}}_{\varepsilon},\varphi_{\varepsilon},\mu_{\varepsilon},\psi_{\varepsilon},\theta_{\varepsilon})$ in the sense of Definition 2.3 with
$\xi_{\varepsilon}:={\beta_{\varepsilon}}(\varphi_{\varepsilon})$ $\in L^{\infty}(0,T;V),$
$\xi_{\Gamma,\varepsilon}:=\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$ $\in L^{\infty}(0,T;V_{\Gamma}).$
Moreover, if the domain $\Omega$ is of class $C^{2}$, it additionally holds
$(\varphi_{\varepsilon},\psi_{\varepsilon})$ $\in L^{2}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)},$
and the equations (2.30)–(2.32) are fulfilled in the strong sense by $\varphi_{\varepsilon}$, $\xi_{\varepsilon}$, $\mu_{\varepsilon}$, $\psi_{\varepsilon}$, $\xi_{\Gamma,\varepsilon}$, and $\theta_{\varepsilon}$.
4.2   Uniform estimates
This section is devoted to derive estimates, uniform with respect to $\varepsilon$, on the approximate solutions $({\boldsymbol{v}}_{\varepsilon},\varphi_{\varepsilon},\xi_{\varepsilon},\mu_{\varepsilon},\psi_{\varepsilon},\xi_{\Gamma,\varepsilon},\theta_{\varepsilon})$. Those will be a key point to obtain suitable convergence properties that allow us to pass to the limit as $\varepsilon\to 0$ later on. In the following, the letter $C$ will denote generic positive constants that may depend on $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3), but not on $\varepsilon$. These constants may also change their value from line to line.
First estimate. To begin with, we test (2.13b) by $\frac{1}{|\Omega|}$ and (2.13c) by $\frac{1}{|\Gamma|}$ to infer that mass conservation for both $\varphi_{\varepsilon}$ and $\psi_{\varepsilon}$ holds as claimed in (1.6)–(1.7). Recalling (2.1) and (2.8) we have
$\mathopen{\langle}\varphi_{\varepsilon}(t)\mathclose{\rangle}_{\Omega}=\mathopen{\langle}\varphi_{0}\mathclose{\rangle}_{\Omega}=m_{0},\quad\mathopen{\langle}\psi_{\varepsilon}(t)\mathclose{\rangle}_{\Gamma}=\mathopen{\langle}\psi_{0}\mathclose{\rangle}_{\Gamma}=:m_{\Gamma 0}\quad\hbox{ for all }\,t\in[0,T].$ (4.5)
This property is intrinsically independent of $\varepsilon$.
We now consider the weak energy dissipation law, already proved in the cases of regular potentials, to $({\boldsymbol{v}}_{\varepsilon},\varphi_{\varepsilon},\mu_{\varepsilon},\psi_{\varepsilon},\theta_{\varepsilon})$, which reads as
$\frac{1}{2}\left\|\nabla\varphi_{\varepsilon}(t)\right\|^{2}_{{\bf H}}+\int_{\Omega}F_{\varepsilon}(\varphi_{\varepsilon}(t))+\frac{1}{2}\left\|\nabla_{\Gamma}\psi_{\varepsilon}(t)\right\|^{2}_{\mathbf{H}_{\Gamma}}$
$\qquad+\int_{\Gamma}G_{\varepsilon}(\psi_{\varepsilon}(t))+\frac{\sigma(K)}{2}\left\|(\psi_{\varepsilon}-\varphi_{\varepsilon})(t)\right\|^{2}_{{H}_{\Gamma}}$
$\qquad+2\int_{0}^{t}\!\!\int_{\Omega}\nu(\varphi_{\varepsilon})|\mathbf{D}{\boldsymbol{v}}_{\varepsilon}|^{2}+\int_{0}^{t}\!\!\int_{\Omega}\lambda(\varphi_{\varepsilon})|{\boldsymbol{v}}_{\varepsilon}|^{2}+\int_{0}^{t}\!\!\int_{\Gamma}\gamma(\psi_{\varepsilon})|{\boldsymbol{v}}_{\varepsilon}|^{2}$
$\qquad+\int_{0}^{t}\!\!\int_{\Omega}M_{\Omega}(\varphi_{\varepsilon})|\nabla\mu_{\varepsilon}|^{2}+\int_{0}^{t}\!\!\int_{\Gamma}M_{\Gamma}(\psi_{\varepsilon})|\nabla_{\Gamma}\theta_{\varepsilon}|^{2}$
$\quad\leq\frac{1}{2}\left\|\nabla\varphi_{0}\right\|^{2}_{{{\bf H}}}+\int_{\Omega}F_{\varepsilon}(\varphi_{0})+\frac{1}{2}\left\|\nabla_{\Gamma}\psi_{0}\right\|^{2}_{\mathbf{H}_{\Gamma}}+\int_{\Gamma}G_{\varepsilon}(\psi_{0})+\frac{\sigma(K)}{2}\left\|\psi_{0}-\varphi_{0}\right\|^{2}_{{H}_{\Gamma}}$ (4.6)
for all $t\in[0,T]$. Now, observe that
$\frac{1}{2}\left\|\nabla\varphi_{0}\right\|^{2}_{{\bf H}}+\int_{\Omega}F_{\varepsilon}(\varphi_{0})+\frac{1}{2}\left\|\nabla_{\Gamma}\psi_{0}\right\|^{2}_{\mathbf{H}_{\Gamma}}+\int_{\Gamma}G_{\varepsilon}(\psi_{0})+\frac{\sigma(K)}{2}\left\|\psi_{0}-\varphi_{0}\right\|^{2}_{{H}_{\Gamma}}\leq C$ (4.7)
since $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ satisfies (2.8) and (4.2a) holds. Hence, in view of (A2) and (A3) and thanks to (4.5) and the Poincaré inequality, it is not difficult to infer that
$\left\|\varphi_{\varepsilon}\right\|_{L^{\infty}(0,T;V)}+\|F_{\varepsilon}(\varphi_{\varepsilon})\|_{L^{\infty}(0,T;L^{1}(\Omega))}+\left\|\psi_{\varepsilon}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}+\|G_{\varepsilon}(\psi_{\varepsilon})\|_{L^{\infty}(0,T;L^{1}(\Gamma))}$
$\quad+\left\|{\boldsymbol{v}}_{\varepsilon}\right\|_{L^{2}(0,T;{\bf V})\cap L^{2}(0,T;{\bf H}_{\Gamma})}+\left\|\nabla\mu_{\varepsilon}\right\|_{L^{2}(0,T;{\bf{H}})}+\left\|\nabla_{\Gamma}\theta_{\varepsilon}\right\|_{L^{2}(0,T;{\bf{H}}_{\Gamma})}\leq C.$ (4.8)
Second estimate. We proceed as in the derivation of (3.13) and (3.15) in the proof of Theorem 2.4. Indeed, let us take an arbitrary test function $\zeta\in L^{2}(0,T;V)$ in (2.13b), then integrate over time and use Hölder’s inequality to obtain that
$\Big{|}\int_{0}^{T}\mathopen{\langle}\partial_{t}\varphi_{\varepsilon},\zeta\mathclose{\rangle}_{V}\Big{|}$ $\leq C\int_{0}^{T}(\left\|\varphi_{\varepsilon}\right\|_{L^{4}(\Omega)}\left\|{\boldsymbol{v}}_{\varepsilon}\right\|_{{\bf{L}}^{4}(\Omega)}+M_{2}\left\|\nabla\mu_{\varepsilon}\right\|_{{\bf H}})\left\|\nabla\zeta\right\|_{{\bf H}}$
$\leq C\left(\left\|\varphi_{\varepsilon}\right\|_{L^{\infty}(0,T;V)}\left\|{\boldsymbol{v}}_{\varepsilon}\right\|_{L^{2}(0,T;{\bf V})}+\left\|\nabla\mu_{\varepsilon}\right\|_{L^{2}(0,T;{\bf H})}\right)\left\|\zeta\right\|_{L^{2}(0,T;V)}$
$\leq C\left\|\zeta\right\|_{L^{2}(0,T;V)}.$
Taking the supremum over all $\zeta\in L^{2}(0,T;V)$ with $\left\|\zeta\right\|_{L^{2}(0,T;V)}\leq 1$, we infer
$\left\|\partial_{t}\varphi_{\varepsilon}\right\|_{L^{2}(0,T;{V^{*}})}\leq C.$ (4.9)
The same argument, acting on equation (2.13c), leads us to infer as well that
$\left\|\partial_{t}\psi_{\varepsilon}\right\|_{L^{2}(0,T;V^{*}_{\Gamma})}\leq C.$ (4.10)
Third estimate. To handle the cases $K>0$ and $K=0$ simultaneously, we introduce the following notation:
$\qquad\alpha(K):=\begin{cases}0&\text{if $K>0$},\\ 1&\text{if $K=0$}.\end{cases}$ (4.11)
We now test (2.13d) by the pair
$(\eta,{\eta_{\Gamma}})=\begin{cases}(\varphi_{\varepsilon}-m_{0},\psi_{\varepsilon}-m_{0})&\text{if $K=0$},\\ (\varphi_{\varepsilon}-m_{0},\psi_{\varepsilon}-m_{\Gamma 0})&\text{if $K>0$},\end{cases}$
which clearly belongs to ${\cal V}_{K}$. After some rearrangements, as well as adding and subtracting the constant $m_{\Gamma 0}$ multiple times in the case $K=0$, we deduce
$\left\|\nabla\varphi_{\varepsilon}\right\|^{2}_{{\bf H}}+\int_{\Omega}\beta_{\varepsilon}(\varphi_{\varepsilon})(\varphi_{\varepsilon}-m_{0})+\left\|\nabla_{\Gamma}\psi_{\varepsilon}\right\|^{2}_{{\bf H}_{\Gamma}}+\int_{\Gamma}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})(\psi_{\varepsilon}-m_{\Gamma 0})$
$\quad=\int_{\Omega}(\mu_{\varepsilon}-\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega})(\varphi_{\varepsilon}-m_{0})+\int_{\Gamma}(\theta_{\varepsilon}-\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma})(\psi_{\varepsilon}-m_{\Gamma 0})$
$\qquad+\sigma(K)\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\big{(}\varphi_{\varepsilon}-\psi_{\varepsilon}-(m_{0}-m_{\Gamma 0})\big{)}$
$\qquad-\int_{\Omega}\pi(\varphi_{\varepsilon})(\varphi_{\varepsilon}-m_{0})-\int_{\Gamma}\pi_{\Gamma}(\psi_{\varepsilon})(\psi_{\varepsilon}-m_{\Gamma 0})$
$\qquad{}+\alpha(K)\int_{\Gamma}(G^{\prime}_{\varepsilon}(\psi_{\varepsilon})-\theta_{\varepsilon})(m_{0}-m_{\Gamma 0}).$ (4.12)
Note that the subtracted mean values $\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega}$ and $\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma}$ in the first two summands on the right-hand side of (4.2) do not change the values of these integrals since, due to (4.5), we have $\mathopen{\langle}\varphi_{\varepsilon}-m_{0}\mathclose{\rangle}_{\Omega}=0$ and $\mathopen{\langle}\psi_{\varepsilon}-m_{\Gamma 0}\mathclose{\rangle}_{\Gamma}=0$.
To deal with the terms on the left-hand side of (4.2), we recall that due to assumption (2.8), $m_{0}$ and $m_{\Gamma 0}$ lie in the interior of the domains $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. We can thus exploit a useful property (see, e.g., [54, Appendix, Prop. A.1] and/or the detailed proof given in [36, p. 908]), namely there exist positive constants $c_{1},c_{2}$ and a nonnegative constant $c_{3}$ such that
$c_{1}\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{1}(\Omega)}+c_{2}\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|_{L^{1}(\Gamma)}-c_{3}$
$\quad\leq\int_{\Omega}\beta_{\varepsilon}(\varphi_{\varepsilon})(\varphi_{\varepsilon}-m_{0})+\int_{\Gamma}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})(\psi_{\varepsilon}-m_{\Gamma 0}).$ (4.13)
For the integrals in the second line of (4.2), we employ Poincaré’s inequality in the bulk and on the surface to obtain that
$\int_{\Omega}(\mu_{\varepsilon}-\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega})(\varphi_{\varepsilon}-m_{0})+\int_{\Gamma}(\theta_{\varepsilon}-\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma})(\psi_{\varepsilon}-m_{\Gamma 0})$
$\quad\leq C\Bigl{(}\left\|\nabla\mu_{\varepsilon}\right\|_{{\bf H}}\left\|\nabla\varphi_{\varepsilon}\right\|_{{\bf H}}+\left\|\nabla_{\Gamma}\theta_{\varepsilon}\right\|_{{\bf H}_{\Gamma}}\left\|\nabla_{\Gamma}\psi_{\varepsilon}\right\|_{{\bf H}_{\Gamma}}\Bigr{)}.$ (4.14)
Moreover, integrals in the third and the fourth line of (4.2) can be bounded by virtue of estimate (4.2) as well as the Lipschitz continuity of $\pi$ and $\pi_{\Gamma}$, so that
$\sigma(K)\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\big{(}\varphi_{\varepsilon}-\psi_{\varepsilon}-(m_{0}-m_{\Gamma 0})\big{)}$
$\quad-\int_{\Omega}\pi(\varphi_{\varepsilon})(\varphi_{\varepsilon}-m_{0})-\int_{\Gamma}\pi_{\Gamma}(\psi_{\varepsilon})(\psi_{\varepsilon}-m_{\Gamma 0})\leq C\bigl{(}\left\|\varphi_{\varepsilon}\right\|^{2}_{H}+\left\|\psi_{\varepsilon}\right\|^{2}_{{H}_{\Gamma}}+1\bigr{)}.$
It remains to estimate the integral in the last line of (4.2), which is only present in the case $K=0$. Recall that if $K=0$, we assumed $\Omega$ to be of class $C^{2}$. Hence, we know from Corollary 4.1 that $(\varphi_{\varepsilon},\psi_{\varepsilon})\in L^{2}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)}$, and that the equations
[…]
$\theta_{\varepsilon}$ $=-\Delta_{\Gamma}\psi_{\varepsilon}-G^{\prime}_{\varepsilon}(\psi_{\varepsilon})+\partial_{\bf n}\varphi_{\varepsilon}$ a.e. on $\Sigma$ (4.16)
hold in the strong sense. Then, with the help of (4.16) and a simple integration by parts, it is not difficult to conclude that
$\alpha(K)\int_{\Gamma}(G^{\prime}_{\varepsilon}(\psi_{\varepsilon})-\theta_{\varepsilon})(m_{0}-m_{\Gamma 0})$
$\quad{}=-\alpha(K)\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}(m_{0}-m_{\Gamma 0})\leq C\,\alpha(K)\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}.$ (4.17)
In the following, we write $\Phi_{\varepsilon}$ to denote generic nonnegative functions
$t\mapsto\Phi_{\varepsilon}(t)\in L^{2}(0,T)\qquad\text{with}\qquad\left\|\Phi_{\varepsilon}\right\|_{L^{2}(0,T)}\leq C\quad\text{for all $\varepsilon>0$}$ (4.18)
i.e., the $L^{2}$-norm is bounded uniformly in $\varepsilon$. Here, “generic” means that the explicit definition of the function $\Phi_{\varepsilon}$ may vary throughout this proof.
All in all, collecting the inequlities (4.2)–(4.2) and (4.17), we conclude that
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon}(t))\right\|_{L^{1}(\Omega)}+\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}(t))\right\|_{L^{1}(\Gamma)}\leq\Phi_{\varepsilon}(t)+C\,\alpha(K)\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}$ (4.19)
for almost all $t\in(0,T)$. Having shown (4.19), now we aim to prove additional $L^{2}$-bounds for the terms $\beta_{\varepsilon}(\varphi_{\varepsilon})$ and $\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$. For that, we take advantage of the growth condition (4.1), which follows from (2.26) in (S2). However, the related analysis has to be performed differently for the cases $K>0$ and $K=0$.
Further estimate in the case $K>0$. As $\alpha(K)=0$ in this case, (4.19) yields
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Omega))}+\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Gamma))}\leq C.$ (4.20)
Of course, thanks to (4.2) we also have
$\left\|F^{\prime}_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Omega))}+\left\|G^{\prime}_{\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Gamma))}\leq C,$
since $\pi$ and $\pi_{\Gamma}$ are Lipschitz continuous. Consequently, by testing (2.13d) first by $(1,0)$ and then by $(0,1)$, one easily realizes that
$\left\|\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega}\right\|_{L^{2}(0,T)}+\left\|\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma}\right\|_{L^{2}(0,T)}\leq C,$ (4.21)
whence, using (4.2) and Poincaré’s inequality once again, we infer that
$\left\|\mu_{\varepsilon}\right\|_{L^{2}(0,T;V)}+\left\|\theta_{\varepsilon}\right\|_{L^{2}(0,T;{V}_{\Gamma})}\leq C.$ (4.22)
Next, recalling that $\sigma(K)=\frac{1}{K}$, we test (2.13d) by $(0,\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}))$ obtaining
$\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}+\int_{\Gamma}\beta^{\prime}_{\Gamma,\varepsilon}(\psi_{\varepsilon})|\nabla_{\Gamma}\psi_{\varepsilon}|^{2}$
$\quad=\int_{\Gamma}(\theta_{\varepsilon}-\pi_{\Gamma}(\psi_{\varepsilon}))\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\frac{1}{K}\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}).$
Observe now that the second term on the left-hand side is nonnegative due to the monotonicity of $\beta_{\Gamma,\varepsilon}$. For the terms on the right-hand side, we use Hölder’s inequality, Young’s inequality and the trace theorem to infer that
$\int_{\Gamma}(\theta_{\varepsilon}-\pi_{\Gamma}(\psi_{\varepsilon}))\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\frac{1}{K}\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}).$
$\quad\leq\frac{1}{2}\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}+C\bigl{(}\left\|\theta_{\varepsilon}\right\|^{2}_{{H}_{\Gamma}}+\left\|\psi_{\varepsilon}\right\|^{2}_{{H}_{\Gamma}}+\left\|\varphi_{\varepsilon}\right\|^{2}_{V}+1\bigr{)}.$
Hence, rearranging the terms and integrating over time we conclude that
$\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;{H}_{\Gamma})}\leq C.$ (4.23)
Next, proceeding similarly, we test (2.13d) by $(\beta_{\varepsilon}(\varphi_{\varepsilon}),0)$. This leads us to
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|^{2}_{H}+\int_{\Omega}\beta^{\prime}_{\varepsilon}(\varphi_{\varepsilon})|\nabla\varphi_{\varepsilon}|^{2}=\int_{\Omega}(\mu_{\varepsilon}-\pi(\varphi_{\varepsilon}))\beta_{\varepsilon}(\varphi_{\varepsilon})+\frac{1}{K}\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\beta_{\varepsilon}(\varphi_{\varepsilon}).$
Again, the second term on the left-hand side is nonnegative owing to (S1), whereas the first term on the right can be easily controlled by Young’s inequality as
$\int_{\Omega}(\mu_{\varepsilon}-\pi(\varphi_{\varepsilon}))\beta_{\varepsilon}(\varphi_{\varepsilon})\leq\frac{1}{2}\left\|\beta_{\varepsilon}(\varphi)\right\|^{2}_{H}+C\bigl{(}\left\|\mu_{\varepsilon}\right\|^{2}_{H}+\left\|\varphi_{\varepsilon}\right\|^{2}_{H}+1\bigr{)}.$
Besides, we handle the last term by combining the monotonicity of $\beta_{\varepsilon}$ with the property in (4.1). Namely, it holds that
$\frac{1}{K}\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\beta_{\varepsilon}(\varphi_{\varepsilon})$
$\quad=-\frac{1}{K}\int_{\Gamma}(\varphi_{\varepsilon}-\psi_{\varepsilon})(\beta_{\varepsilon}(\varphi_{\varepsilon})-\beta_{\varepsilon}(\psi_{\varepsilon}))+\frac{1}{K}\int_{\Gamma}(\psi_{\varepsilon}-\varphi_{\varepsilon})\beta_{\varepsilon}(\psi_{\varepsilon})$
$\quad\leq\frac{1}{K}\int_{\Gamma}|\psi_{\varepsilon}-\varphi_{\varepsilon}|\,|\beta_{\varepsilon}(\psi_{\varepsilon})|$
$\quad\leq\frac{\kappa_{1}}{K}\int_{\Gamma}(|\psi_{\varepsilon}|+|\varphi_{\varepsilon}|)|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})|+\frac{\kappa_{2}}{K}\int_{\Gamma}(|\psi_{\varepsilon}|+|\varphi_{\varepsilon}|)$
$\quad\leq\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}+C\bigl{(}\left\|\psi_{\varepsilon}\right\|^{2}_{{H}_{\Gamma}}+\left\|\varphi_{\varepsilon}\right\|^{2}_{V}+1\bigr{)}.$
Hence, with the help of (4.23), this shows the corresponding estimate
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;H)}\leq C.$ (4.24)
Further estimate in the case $K=0$. Recall that, as $K=0$, it now holds that $\sigma(K)=0$, $\alpha(K)=1$, and $\varphi_{\varepsilon}|_{\Gamma}=\psi_{\varepsilon}$ a.e. on $\Sigma$, along with (4.15) and (4.16). Here, in our argumentation, we follow in parts the procedure devised in [19].
Multiplying (4.15) by $1/|\Omega|$ and integrating over $\Omega$, we find that
$|\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega}|\leq C\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}+\frac{1}{|\Omega|}\bigl{(}\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{1}(\Omega)}+\left\|\pi(\varphi_{\varepsilon})\right\|_{L^{1}(\Omega)}\bigr{)}.$ (4.25)
Similarly, multiplying (4.16) by $1/|\Gamma|$ and integrating over $\Gamma$, we infer that
$|\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma}|\leq C\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}+\frac{1}{|\Gamma|}\bigl{(}\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|_{L^{1}(\Gamma)}+\left\|\pi_{\Gamma}(\psi_{\varepsilon})\right\|_{L^{1}(\Gamma)}\bigr{)}.$ (4.26)
Then, combining (4.25) and (4.26), on account of the estimates (4.2) and (4.19) along with the Lipschitz continuity of $\pi$ and $\pi_{\Gamma}$, we deduce that
$|\mathopen{\langle}\mu_{\varepsilon}\mathclose{\rangle}_{\Omega}|+|\mathopen{\langle}\theta_{\varepsilon}\mathclose{\rangle}_{\Gamma}|\leq C\bigl{(}\Phi_{\varepsilon}+\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}\bigr{)}.$ (4.27)
Combining (4.2) and (4.7), we obtain the estimate
$\left\|\nabla\mu_{\varepsilon}\right\|_{{{\bf{H}}}}+\left\|\nabla_{\Gamma}\theta_{\varepsilon}\right\|_{{{\bf{H}}_{\Gamma}}}\leq\Phi_{\varepsilon}.$
Hence, with the help of Poincaré’s inequality, we arrive at
$\left\|\mu_{\varepsilon}(t)\right\|_{V}+\left\|\theta_{\varepsilon}(t)\right\|_{V_{\Gamma}}\leq C\bigl{(}\Phi_{\varepsilon}(t)+\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}\bigr{)}$ (4.28)
for almost all $t\in(0,T)$. Now, we multiply (4.15) by $\beta_{\varepsilon}(\varphi_{\varepsilon})$ and integrate by parts. This yields
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|^{2}_{H}+\int_{\Omega}\beta^{\prime}_{\varepsilon}(\varphi_{\varepsilon})|\nabla\varphi_{\varepsilon}|^{2}$
$\quad=\int_{\Omega}(\mu_{\varepsilon}-\pi(\varphi_{\varepsilon}))\beta_{\varepsilon}(\varphi_{\varepsilon})+\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\varepsilon}(\varphi_{\varepsilon}).$
$\quad\leq\frac{1}{2}\int_{\Omega}|\mu_{\varepsilon}-\pi(\varphi_{\varepsilon}))|^{2}+\frac{1}{2}\left\|\beta_{\varepsilon}(\varphi_{\varepsilon})\right\|^{2}_{H}+\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\varepsilon}(\varphi_{\varepsilon}).$ (4.29)
Similarly, multiplying (4.16) by $-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$, it is straightforward to deduce that
$\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}+\int_{\Gamma}\beta^{\prime}_{\Gamma,\varepsilon}(\psi_{\varepsilon})|\nabla_{\Gamma}\psi_{\varepsilon}|^{2}$
$\quad=\int_{\Gamma}(\theta_{\varepsilon}-\pi_{\Gamma}(\psi_{\varepsilon}))\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$
$\quad\leq\int_{\Omega}|\theta_{\varepsilon}-\pi_{\Gamma}(\psi_{\varepsilon})|^{2}+\frac{1}{4}\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}-\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}).$ (4.30)
Recalling (4.1), we observe that
$\biggl{|}\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\varepsilon}(\varphi_{\varepsilon})-\int_{\Gamma}\partial_{\bf n}\varphi_{\varepsilon}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\biggr{|}$
$\quad{}\leq\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}\left\|(\kappa_{1}+1)|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})|+\kappa_{2}\right\|_{{H}_{\Gamma}}$
$\quad{}\leq\frac{1}{4}\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|^{2}_{{H}_{\Gamma}}+C\bigl{(}\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{{H}_{\Gamma}}^{2}+1\bigr{)}.$
Hence, adding (4.2) and (4.2), and using (4.2) as well as (4.28), we conclude that
$\left\|\beta_{\varepsilon}(\varphi_{\varepsilon}(t))\right\|_{H}+\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon}(t))\right\|_{{H}_{\Gamma}}\leq C\bigl{(}\Phi_{\varepsilon}(t)+\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}\bigr{)}$ (4.31)
for almost all $t\in(0,T)$. Now, recalling again (4.15) and (4.16), we observe that $\varphi_{\varepsilon}$ solves the following bulk-surface elliptic problem:
$-\Delta\varphi_{\varepsilon}$ $=\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})$ in $\quad\text{in $\Omega$},$
$-\Delta_{\Gamma}\psi_{\varepsilon}+\partial_{\bf n}\varphi_{\varepsilon}$ $=\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\pi_{\Gamma}(\psi_{\varepsilon})$ on $\quad\text{on $\Gamma$},$
$\varphi_{\varepsilon}|_{\Gamma}$ $=\psi_{\varepsilon}$ on $\Gamma$
a.e. in $(0,T)$. Due to (4.2), (4.28), (4.31) and the Lipschitz continuity of $\pi$ and $\pi_{\Gamma}$, it is clear that the right-hand sides in the above system belong to $L^{2}(\Omega)$ and $L^{2}(\Gamma)$, respectively. Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we deduce that the estimate
$\left\|\varphi_{\varepsilon}\right\|_{H^{2}(\Omega)}+\left\|\psi_{\varepsilon}\right\|_{H^{2}(\Gamma)}$
$\quad{}\leq C\Bigl{(}\left\|\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})\right\|_{H}+\left\|\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})+\psi_{\varepsilon}-\pi_{\gamma}(\psi_{\varepsilon})\right\|_{{H}_{\Gamma}}\Bigl{)}$
holds a.e. in $(0,T)$. Now, in view of (4.2), (4.28), (4.31) we can completely control the above right-hand side and infer that
$\left\|\varphi_{\varepsilon}(t)\right\|_{H^{2}(\Omega)}+\left\|\psi_{\varepsilon}(t)\right\|_{H^{2}(\Gamma)}\leq C\bigl{(}\Phi_{\varepsilon}(t)+\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}\bigr{)}$ (4.32)
for almost all $t\in(0,T)$. On this basis, at this point we can use the standard trace theorem for the normal derivative concluding that, for some fixed 3/$23/2<s<$ 2 there is a positive constant $C_{s}$ such that
$\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}\leq C_{s}\left\|\varphi_{\varepsilon}(t)\right\|_{H^{s}(\Omega)}$
for almost all $t\in(0,T)$. Hence, as $H^{2}(\Omega)\subset H^{s}(\Omega)\subset V$ with compact embeddings, we infer from (4.32) by means of the Ehrling lemma that
$\left\|\varphi_{\varepsilon}(t)\right\|_{H^{2}(\Omega)}+\left\|\psi_{\varepsilon}(t)\right\|_{H^{2}(\Gamma)}+\left\|\partial_{\bf n}\varphi_{\varepsilon}(t)\right\|_{{H}_{\Gamma}}$
$\quad\leq C\bigl{(}\Phi_{\varepsilon}(t)+C_{s}\left\|\varphi_{\varepsilon}(t)\right\|_{H^{s}(\Omega)}\bigr{)}+C_{s}\left\|\varphi_{\varepsilon}(t)\right\|_{H^{s}(\Omega)}$
$\quad\leq\delta\left\|\varphi_{\varepsilon}(t)\right\|_{H^{2}(\Omega)}+C\,\Phi_{\varepsilon}(t)+C\delta^{-1}\left\|\varphi_{\varepsilon}(t)\right\|_{V},$ (4.33)
for all $t\in(0,T)$ and any $\delta\in(0,1)$. Eventually, from (4.2) (4.2), it follows that
$\left\|\varphi_{\varepsilon}\right\|_{L^{2}(0,T;H^{2}(\Omega))}+\left\|\psi_{\varepsilon}\right\|_{L^{2}(0,T;H^{2}(\Gamma))}+\left\|\partial_{\bf n}\varphi_{\varepsilon}\right\|_{L^{2}(0,T;{H}_{\Gamma})}\leq C$ (4.34)
and consequently, recalling (4.28) and (4.31), we also have
$\left\|\mu_{\varepsilon}\right\|_{L^{2}(0,T;V)}+\left\|\theta_{\varepsilon}\right\|_{L^{2}(0,T;V_{\Gamma})}+\left\|\beta_{\varepsilon}(\varphi_{\varepsilon}(t))\right\|_{L^{2}(0,T;H)}+\left\|\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;{H}_{\Gamma})}\leq C.$ (4.35)
4.3   Passage to the limit and conclusion of the proof
The final step consists in passing to the limit as $\varepsilon$ is sent to zero. As the line of argument resembles the one presented in Section 3.1.4, we proceed rather quickly just pointing out the main points and differences.
Owing to the above uniform estimates and to standard compactness results, we obtain that there exist a subsequence of $\varepsilon$ and a seventuple of limits
$({\boldsymbol{v}}^{*},\varphi^{*},\xi^{*},\mu^{*},\psi^{*},\xi_{\Gamma}^{*},\theta^{*})$
such that, as $\varepsilon\to 0$,
${\boldsymbol{v}}_{\varepsilon}$ $\to{\boldsymbol{v}}^{*}\quad$ weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$,
strongly in $C^{0}([0,T];H^{s}(\Omega))$, and a.e. in $Q$,
${\boldsymbol{v}}_{\varepsilon}|_{\Gamma}$ $\to{\boldsymbol{v}}^{*}|_{\Gamma}\quad$ weakly in $L^{2}(0,T;{\bf H}_{\Gamma})$,
strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and a.e. on $\Sigma$,
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\quad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\quad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$
${\beta_{\varepsilon}}(\varphi_{\varepsilon})$ $\to\xi^{*}\quad$ weakly in $L^{2}(0,T;{H})$,
$\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$ $\to\xi_{\Gamma}^{*}\quad$ weakly in $L^{2}(0,T;{H_{\Gamma}})$,
$\mu_{\varepsilon}$ $\to\mu^{*}\quad$ weakly in $L^{2}(0,T;V)$,
$\theta_{\varepsilon}$ $\to\theta^{*}\quad$ weakly in $L^{2}(0,T;{V}_{\Gamma})$,
for all $s\in[0,1)$. In the case $K=0$, we further infer from (4.34) the convergences
$\varphi_{\varepsilon}$ $\to\varphi^{*}$ weakly in $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Omega))$},$
$\psi_{\varepsilon}$ $\to\psi^{*}$ weakly in $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Gamma))$}.$
Repeating the arguments employed in Section 3.1.4, we can easily show that the above weak and strong convergences suffice to pass to the limit in the variational formulation (2.28a)–(2.28d) written for $\beta={\beta_{\varepsilon}}$ and ${\beta_{\Gamma}}={\beta_{\Gamma,\varepsilon}}$. Furthermore, the inclusions
$\xi^{*}\in\beta(\varphi^{*})$ a.e. in $Q$  and  $\xi_{\Gamma}^{*}\in{\beta_{\Gamma}}(\psi^{*})$ a.e. on $\Sigma$
follow directly from the maximality of the monotone operators $\beta$ and $\beta_{\Gamma}$, and the facts that
$\lim_{\varepsilon\to 0}\int_{0}^{T}\!\!\int_{\Omega}\beta_{\varepsilon}(\varphi_{\varepsilon})\varphi_{\varepsilon}=\int_{0}^{T}\!\!\int_{\Omega}\xi^{*}\varphi^{*},\quad\lim_{\varepsilon\to 0}\int_{0}^{T}\!\!\int_{\Gamma}\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})\psi_{\varepsilon}=\int_{0}^{T}\!\!\int_{\Gamma}\xi_{\Gamma}^{*}\psi^{*}$
(see, e.g., [9, Prop. 1.1, p. 42]). Due to the aforementioned strong convergences of $\varphi_{\varepsilon}$ and $\psi_{\varepsilon}$, it is straightforward to check that condition (iii) of Definition 2.8 is fulfilled. Moreover, condition (iv) of Definition 2.8 can be established by proceeding analogously as in Subection 3.1.4.
Finally, if the domain is of class $C^{2}$, we need to establish the higher regularity properties of the phase-fields $\varphi_{*}$ and $\psi_{*}$. In the case $K=0$, this directly follows from the above convergences. In the case $K>0$, these properties can be proved as in Subsection 3.1.5 by taking advantage of the regularities $L^{2}(0,T;{H})$ for $\xi$ and $L^{2}(0,T;{H_{\Gamma}})$ for $\xi_{\Gamma}$. This concludes the proof of Theorem 2.9.
Appendix: Some calculus for bulk-surface function spaces
Proposition A.1.
Let $T>0$ and $K\geq 0$ be arbitrary.
  1. (a)
    Let $(u,v)\in L^{2}(0,T;{\cal V}_{K})$ and suppose that the weak time derivative satisfies $(\partial_{t}u,\partial_{t}v)\in L^{2}(0,T;{\cal V}_{K}^{*})$. Then, the continuity property $(u,v)\in C^{0}([0,T];{\cal H})$ holds, the mapping
    $t\mapsto\left\|(u,v)(t)\right\|_{{\cal H}}^{2}=\left\|u(t)\right\|_{H}^{2}+\left\|v(t)\right\|_{{H}_{\Gamma}}^{2}$
    is absolutely continuous in $[0,T]$, and the chain rule formula
    $\frac{\mathrm{d}}{\mathrm{d}t}\Big{[}\left\|u(t)\right\|_{H}^{2}+\left\|v(t)\right\|_{{H}_{\Gamma}}^{2}\Big{]}=2\big{<}(\partial_{t}u,\partial_{t}v)(t),(u,v)(t)\big{>}_{{\cal V}_{K}}$ (A.1)
    holds for almost all $t\in[0,T]$.
  2. (b)
    Let $(u,v)\in L^{2}\big{(}0,T;H^{3}(\Omega)\times H^{3}(\Gamma)\big{)}$ with $K\partial_{n}u=v-u$ a.e. on $\Sigma$, and suppose that their weak time derivative satisfies $(\partial_{t}u,\partial_{t}v)\in L^{2}(0,T;{{\cal V}}^{*})$. Then, the continuity property $(u,v)\in C^{0}([0,T];{\cal V}_{K})$ holds, the mapping
    $t\mapsto\left\|\nabla u(t)\right\|_{\bf{H}}^{2}+\left\|\nabla_{\Gamma}v(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v(t)-u(t)\right\|^{2}_{{H}_{\Gamma}}$
    is absolutely continuous in $[0,T]$, and the chain rule formula
    $\frac{\mathrm{d}}{\mathrm{d}t}\Big{[}\left\|\nabla u(t)\right\|_{\bf{H}}^{2}+\left\|\nabla_{\Gamma}v(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v(t)-u(t)\right\|^{2}_{{H}_{\Gamma}}\Big{]}$
    $\quad=2\big{<}(\partial_{t}u,\partial_{t}v)(t),(-\Delta u,-\Delta_{\Gamma}v+\partial_{\bf n}u)(t)\big{>}_{{\cal V}}$ (A.2)
    holds for almost all $t\in[0,T]$.
Proof.
Proof of (a). Since ${\cal V}_{K}$ and ${\cal H}$ are separable Hilbert spaces with compact embedding ${\cal V}_{K}{\hookrightarrow}{\cal H}$ and continuous embedding ${\cal H}{\hookrightarrow}{\cal V}^{*}_{K}$, the assertion follows directly from the Lions–Magenes lemma (see, e.g., [60, Chapter III, Lemma 1.2]).
Proof of (b). We first fix $u$ and $v$ as arbitrary representatives of their respective equivalence class. Recall that due to (a), we have $u\in C^{0}([0,T];H)$ and $v\in C^{0}([0,T];{H}_{\Gamma})$. We can thus extend the functions $u$ and $v$ onto $[-T,0]$ by defining $u(t)$ and $v(t)$ by reflection for all $t<0$.
Let $\rho\in C^{\infty}_{c}({\mathbb{R}})$ be a nonnegative function with $\mathop{\rm supp}\nolimits\rho\subset(0,1)$ and $\left\|\rho\right\|_{L^{1}({\mathbb{R}})}=1$. For any $k\in{\mathbb{N}}$, we set
$\rho_{k}(s):=k\rho(ks)\quad\text{for all $s\in{\mathbb{R}}$}.$
For any Banach space $X$ and any function $f\in L^{2}(-1,T;X)$, we define
$f_{k}(t):=(\rho_{k}*f)(t)=\int_{t-\frac{1}{k}}^{t}\rho_{k}(t-s)\,f(s)\,{\mathrm{d}s}$
for all $t\in[0,T]$ and all $k\in{\mathbb{N}}$. By this construction, we have $f_{k}\in C^{\infty}([0,T];X)$ with $f_{k}\to f$ strongly in $L^{2}(0,T;X)$ as $k\to\infty$.
For any $k\in{\mathbb{N}}$, we now choose $X=H^{3}(\Omega)$ to define $u_{k}$ and $X=H^{3}(\Gamma)$ to define $v_{k}$ as described above. By this construction, it holds $\partial_{t}u_{k}=(\partial_{t}u)_{k}$ and $\partial_{t}\nabla u_{k}=\nabla\partial_{t}u_{k}$ a.e. in $Q$ as well as $\partial_{t}v_{k}=(\partial_{t}v)_{k}$ and $\partial_{t}\nabla_{\Gamma}v_{k}=\nabla_{\Gamma}\partial_{t}v_{k}$ a.e. on $\Sigma$ for all $k\in{\mathbb{N}}$. Moreover, as $k\to\infty$, we have
$u_{k}$ $\to u$ strongly in $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Omega))$},$ (A.3)
$v_{k}$ $\to v$ strongly in $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Gamma))$},$ (A.4)
$(u_{k},v_{k})$ $\to(u,v)$ strongly in $\quad\text{strongly in $L^{2}(0,T;{\cal V}_{K})$},$ (A.5)
$(\partial_{t}u_{k},\partial_{t}v_{k})$ $\to(\partial_{t}u,\partial_{t}v)$ strongly in $\quad\text{strongly in $L^{2}(0,T;{\cal V}^{*})$}.$ (A.6)
In the following, the letter $C$ will denote generic positive constants that are independent of $k$ and may change their value from line to line. Now, for any $k\in{\mathbb{N}}$, we derive the identity
$\frac{\mathrm{d}}{\mathrm{d}t}\Big{[}\left\|\nabla u_{k}\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v_{k}\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v_{k}-u_{k}\right\|^{2}_{{H}_{\Gamma}}\Big{]}$
$\quad=2\big{<}(\partial_{t}u_{k},\partial_{t}v_{k}),(-\Delta u_{k},-\Delta_{\Gamma}v_{k}+\partial_{\bf n}u_{k})\big{>}_{{{\cal V}}}$ (A.7)
in $[0,T]$ by differentiating under the integral sign, applying integration by parts, and employing the relation
$\sigma(K)(v_{k}-u_{k})=\begin{cases}0&\text{if $K=0$},\\ \partial_{\bf n}u_{k}&\text{if $K>0$},\end{cases}\qquad\hbox{a.e.\ on~{}$\Sigma$}.$
Let now $j,k\in{\mathbb{N}}$ be arbitrary. Proceeding as above, we calculate
$\frac{\mathrm{d}}{\mathrm{d}t}\Big{[}\left\|\nabla u_{j}-\nabla u_{k}\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v_{j}-\nabla_{\Gamma}v_{k}\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|(v_{j}-v_{k})-(u_{j}-u_{k})\right\|^{2}_{{H}_{\Gamma}}\Big{]}$
$\quad=2\big{<}\big{(}\partial_{t}(u_{j}-u_{k}),\partial_{t}(v_{j}-v_{k})\big{)},\big{(}-\Delta(u_{j}-u_{k}),-\Delta_{\Gamma}(v_{j}-v_{k})+\partial_{\bf n}(u_{j}-u_{k})\big{)}\big{>}_{{{\cal V}}}$
$\qquad\leq C\Big{(}\left\|\big{(}\partial_{t}(u_{j}-u_{k}),\partial_{t}(v_{j}-v_{k})\big{)}\right\|_{{\cal V}_{*}}^{2}+\left\|u_{j}-u_{k}\right\|_{H^{3}(\Omega)}^{2}+\left\|v_{j}-v_{k}\right\|_{H^{3}(\Gamma)}^{2}\Big{)}$ (A.8)
in $[0,T]$. Here, we have used the embedding $H^{3}(\Omega){\hookrightarrow}H^{2}(\Gamma)$ resulting from the trace theorem, which yields
$\left\|\partial_{\bf n}(u_{j}-u_{k})\right\|_{{V}_{\Gamma}}\leq\left\|u_{j}-u_{k}\right\|_{H^{2}(\Gamma)}\leq C\left\|u_{j}-u_{k}\right\|_{H^{3}(\Omega)}.$
Let now $s,t\in[0,T]$ be arbitrary with $s\leq t$. We then integrate inequality (A.8) with respect to time from $s$ to $t$. This yields
$\left\|(\nabla u_{j}-\nabla u_{k})(t)\right\|_{{\bf{H}}}^{2}+\left\|(\nabla_{\Gamma}v_{j}-\nabla_{\Gamma}v_{k})(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}$
$\quad+\sigma(K)\left\|(v_{j}(t)-v_{k}(t))-(u_{j}(t)-u_{k}(t))\right\|^{2}_{{H}_{\Gamma}}$
$\leq\left\|(\nabla u_{j}-\nabla u_{k})(s)\right\|_{{\bf{H}}}^{2}+\left\|(\nabla_{\Gamma}v_{j}-\nabla_{\Gamma}v_{k})(s)\right\|_{\mathbf{H}_{\Gamma}}^{2}$ (A.9)
$\quad+\sigma(K)\left\|(v_{j}(s)-v_{k}(s))-(u_{j}(s)-u_{k}(s))\right\|^{2}_{{H}_{\Gamma}}$
$\quad+C\int_{s}^{t}\left\|\big{(}\partial_{t}(u_{j}-u_{k}),\partial_{t}(v_{j}-v_{k})\big{)}\right\|_{{{\cal V}^{*}}}^{2}+\left\|u_{j}-u_{k}\right\|_{H^{3}(\Omega)}^{2}+\left\|v_{j}-v_{k}\right\|_{H^{3}(\Gamma)}^{2}.$
Since $(u_{k},v_{k})\to(u,v)$ strongly in $L^{2}\big{(}0,T;(H^{3}(\Omega)\times H^{3}(\Gamma))\big{)}$, we can fix $s\in[0,t]$ such that $(u_{k},v_{k})(s)\to(u,v)(s)$ strongly in $H^{3}(\Omega)\times H^{3}(\Gamma)$. Recalling the convergences (A.3)–(A.6), we thus infer that the right-hand side in (Proof.) tends to zero as $j,k\to\infty$. Consequently, $(\nabla u_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];{\bf{H}})$ and $(\nabla v_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];\mathbf{H}_{\Gamma})$. We thus conclude
$\nabla u_{k}$ $\to\nabla u$ strongly in $\quad\text{strongly in $C^{0}([0,T];{\bf{H}})$},$ (A.10)
$\nabla_{\Gamma}v_{k}$ $\to\nabla_{\Gamma}u$ strongly in $C^{0}([0,T];\mathbf{H}_{\Gamma})$ (A.11)
as $k\to\infty$. Together with (a), this proves
$(u,v)\in C^{0}([0,T];{\cal V}_{K}).$
Let now $s,t\in[0,T]$ be arbitrary. Without loss of generality, we assume $s\leq t$. Integrating (A.7) with respect to time from $s$ to $t$, we obtain
$\left\|\nabla u_{k}(t)\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v_{k}(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v_{k}(t)-u_{k}(t)\right\|^{2}_{{H}_{\Gamma}}$
$\quad=\left\|\nabla u_{k}(s)\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v_{k}(s)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v_{k}(s)-u_{k}(s)\right\|^{2}_{{H}_{\Gamma}}$
$\qquad+2\int_{s}^{t}\big{<}(\partial_{t}u_{k},\partial_{t}v_{k}),(-\Delta u_{k},-\Delta_{\Gamma}v_{k}+\partial_{\bf n}u_{k})\big{>}_{{{\cal V}}}.$
Invoking the convergences (A.3)–(A.6), (A.10) and (A.11), we pass to the limit $k\to\infty$ in this identity. This yields
$\left\|\nabla u(t)\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v(t)-u(t)\right\|^{2}_{{H}_{\Gamma}}$
$\quad=\left\|\nabla u(s)\right\|_{{\bf{H}}}^{2}+\left\|\nabla_{\Gamma}v(s)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v(s)-u(s)\right\|^{2}_{{H}_{\Gamma}}$
$\qquad+2\int_{s}^{t}\big{<}\big{(}\partial_{t}u,\partial_{t}v\big{)},\big{(}-\Delta u,-\Delta_{\Gamma}v+\partial_{\bf n}u\big{)}\big{>}_{{{\cal V}}}.$
As the integrand of the integral on the right-hand side belongs to $L^{1}(0,T)$, we conclude that the mapping $t\mapsto\left\|\nabla u(t)\right\|_{\bf{H}}^{2}+\left\|\nabla_{\Gamma}v(t)\right\|_{\mathbf{H}_{\Gamma}}^{2}+\sigma(K)\left\|v(t)-u(t)\right\|^{2}_{{H}_{\Gamma}}$ is absolutely continuous in $[0,T]$. It is thus differentiable almost everywhere in $[0,T]$ and its derivative satisfies the formula ((b)). This verifies (b) and thus, the proof is complete. ∎
Acknowledgement
This research has been performed within the framework of the MIUR-PRIN Grant 2020F3NCPX “Mathematics for industry 4.0 (Math4I4)”. The present paper also benefits from the support of the GNAMPA (Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni) of INdAM (Istituto Nazionale di Alta Matematica).
Andrea Signori has been supported by “MUR GRANT Dipartimento di Eccellenza” 2023-2027.
Moreover, Patrik Knopf was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): on the one hand by the DFG-project 524694286, and on the other hand by the RTG 2339 “Interfaces, Complex Structures, and Singular Limits”. Their support is gratefully acknowledged.
Parts of this work were done while Patrik Knopf was visiting the Dipartimento di Matematica “F. Casorati” of the Università di Pavia whose hospitality is gratefully appreciated.
Last but not least, we thank Jonas Stange for helping us to proofread the manuscript.